A Complete Bibliography of Publications in Japanese Studies in the History of Science

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

08 October 2018 Version 1.00

Title word cross-reference

18 × 26 [Yua74]. 43.5 × 31 [Ish75]. α [Nis65].

-Rays [Nis65].

000 [Ish75].

1 [Sat79a]. 1/1962 [Ano69c]. 11th [YL72]. 11th-Century [YL72]. 15th [SW67]. 16th [Yua62]. 18th [Yag66]. 1930 [Tan65]. 1965 [Nak67a]. 1968 [Ano69c, Yaj68]. 1969 [Yab69]. 1970 [TH70]. 19th [Dôk73, Ish75].

20th [Yua62].

36 [Wat70a]. 367 [Ish78].

40pp. [Yam74]. 42pp [Ish73].

560 [Mur75b].

7 [Ano69c]. 7/1968 [Ano69c]. 70 [Aok73].

800 [Ish74].

Abū [YL71]. Abbe [iT73, iT75, Yam76]. Abbot [Yam76]. Abhorrence [Nak64a]. abstract [Miu62, Nak62b, Tom62, Yos62b]. Abu [Lev70]. Academic [WO68]. Acadèmie [Yaj68]. Academy [Ish73]. Accuracy [Nak63]. Achievement [Ish74, Tak76]. Across [Mat78, Wat70b]. Activities [Hig75, Hir70a, Nak65e, Shi65, Shi69]. Activity [Yua62]. actuel [Mur76]. Adolphe [Yos77]. African [FY68]. after [FY68]. Against [Tan66, Yaj63a]. Agassi [Wat70a]. Age [Mat75, Tom62, Tan77a]. Ages [Miy69, Tak76]. Akira [Yua74]. Alchemist [YL72]. Alchemy [Siv67]. Alleged [Nak68]. Alten [Sai62]. Alternative [Sat79b]. am [Kan74]. Amber [BM72]. America [Nak67a]. American [Mat78, Wat66a, Wat70b, Wat76]. Anaesthetic [Miy66]. Analysis [TH75, WO68]. Analytical [Miy77]. Anatomical [Ach66]. Ancient [Iwa77, Mur68, Ots75, Suz65]. Anerkennung [Tan63]. Anethesia [Tak67]. Anniversary [Yag75]. antiken [Ots75]. Aoki [Ano77b, Sip72]. Appreciation [Ish74]. Approach [Nak72, SH78, YK73]. Arabic [Lev71]. Archaeological [Mur70]. Archimedes [Sat79a]. Armin [Hir70b]. Art [Tan65]. Arts [Yos62b]. Asada [Nak68]. Aschendorff [Aok73]. Asia [Pre74]. Asiatic [Bla73]. Aspect [Kob64]. Aspects [Lev71, Sib62, Yam71]. Astrology [Nak62b]. Astronomical [Bla74, Nak64b, Yab63]. Astronomy [Ait75, Nak62b, Nak62d, Nak65e, Yab69]. Atom [HN70, Nis67]. Atomhypothese [Tan66]. Atomic [HN64, Nis65, Nis69, Tan66, Tan77a, Yag64, Yag67, Yag69, Yag72, Yos77]. atomique [Yos77]. Atomism [Fuj75, Oha67, Tan69, Tan62, Tan63, Tan66, Tan72]. Atomistic [Mur67, Tan69]. Atomistik [Tan62, Tan63, Tan66, Tan72, Tan62]. Attempt [Sat79a]. Attitude [Yam77a]. Auffassungen [Tan66]. Aufnahme [Tan67]. avant [Yaj68, Yam77b]. Avogadro [Tan63]. Avogadroschen [Tan63].

Background [Yab69]. Backgrounds [SH78]. Bacra [Lev71]. Baden [Hir70b]. Baifûkan [Ish78]. Ballistical [Ita63]. Balmer [Nis66]. Basic [Tan76, Yam77a]. Bd [Aok73]. be [Shi72]. Bead [Mur68]. Beads [Mur68]. before [Hir66, Ish73, II62, Yam77b]. Beginning [Ima78]. Beginnings [Sip72]. Begriff [Tan72]. Begriffs [Tan72]. Beikokujin [Mat78]. beim [Tan62]. Beitrag [Tan62, Tan63, Tan66, Tan72]. Beiträge [Aok73, Tan76]. Bengal [Bla73]. Benjamin [Yam76]. Beschreibungen [Ots75]. Between [Shi66, Dok67, Mic79]. Binominal [Kim77]. Biochemistry [Dok69]. bioethics [Ots77]. Biography [Kob76, Yag71]. Biology [Nak65b, Suz72, YT62]. Blood [Nak74]. Bohr [HN64, HN70, Nis67]. Book [Aok71, Aok73, Eza77, Got77, Hig75, Hir70b, Ish73, Ish74, Ish75, Ish76, Ish78, Kaw79, Kim74, Mat78, Miu79, Mur78, Mur71, Mur75b, Nak67a, Nak76, Nar76, Shi72, Tak76, Tak77a, Tak77b, TH70, TO78, Wat70a, Wat74, Yab69, Yaj68, Yam73, Yam74, Yua74, Yok72a]. Books [Hir69, Mac74]. Born [Kon78]. Brief [Ish70, Kob76]. Broek [Hir71, Mac77]. Bunmei [Tak77a]. Burger [Ish70].

C [Shi66, Tak76]. C. [Shi66]. Cabinet [Mac75]. calcul [Yam71]. calculating [Yam71]. Calendar [Siv65]. Calendrical [Nak65d]. Cambridge [Aok71, Wat74]. can [Shi72]. Cartesian [Yam76]. cartésienne [Yam76]. Cartography [Ish74]. Cavendish [Mat75]. Ce [Yaj63a]. Centenary [yH77]. Center [Yua62]. Centrifugal [Taj76]. Century [Dôk73, HN68, Ish75, WO68, YL72, Yua62, Yua63, SW67, Tan66, Wat70b, Yag66]. Changes [Yok72a]. Characteristics [Nak65d, Nem74]. Characters [Siv65]. Charles [Yos77]. Chart [Ach66]. Chemical [Lev71, Mur68, Mur70, Nis67, Tan69, Tan76, Tan62, Tan65, Tan72, Hig75]. Chemie [Tan64, Tan65, Tan66, Tan67, Tan76]. Chemiker [Tan76]. Chemischen [Tan62, Tan72]. chemischer [Tan65, Tan76]. Chemist [Tan76]. Chemistry [Hig75, ÔD73, iT69, Tan64, Tan65, Tan66, Tan67, Tan68, TY65, TY68, Yag66, Tan76]. Chhi [Sai62]. Chimique [Mat75]. China [Miy77, Miy66, Nak66, NGD66, Tom62, Yab63, Yos62a]. Chinese [Nak70, Yab69, Miy69, Mur68, Nak65d, Nak69, Siv67, Yok72a, YL72]. Ch’ing [Yab63]. Chronology [Ish70]. Chuokoron [Shi72]. Circle [Sat79a]. Circulation [Nak74]. civilisations [Yaj63a]. Civilizations [Tak77a, Yaj63a]. Classical [Eza79, Tan72]. Classification [SS66]. Classified [TH73]. cm [Ish75, Yua74]. Collected [Yua74]. Colleges [Nar76]. Colonialism [Nak71]. Combination [Nis66]. Committee [Yag71]. Commodities [SS66]. Communities [Yua70]. Community [Nar76]. Comparative [SH78]. Compilation [TH72]. Computer [TH72]. Concept [Hir62a, Taj75, Tan69, iT75, Tan72]. Concepts [Tom62, Tan67]. Conceptual [Tan69]. concerning [Hir65]. Conditions [Hir63]. Condorcet [Yam71]. conduction [Yam77b, Yam77b]. Confucianism [Miy77]. Congress [Pre74]. Consciousness [Has64]. conservation [Yam76]. Consideration [Hir65]. Considerations [Nis67, Taj74, TH75]. Constitution [HN64]. Contacts [Wat70b]. Contemporary [Hsi68, WO68, Mat78]. Contents [Ano69c, Ano70c, Ano72c, Ano73c, Ano74d, Ano76c, Ano77e, Ano78d, Ano78e, Ano79c, Ano79d]. Context [Bla74]. Contribute [Nak62b]. Contribution [Tan69, Tan62, Tan63, Tan66, Tan72]. Contributions [Eza79, Yam62, Tan76]. Contributors [Ano69b, Ano70a, Ano71b, Ano72a, Ano73a, Ano74b, Ano76a, Ano77c, Ano78b, Ano79a]. Controversy [Dok67]. Conversion [Siv65]. Copernicanism [Nak64a]. Copernicus [Yaj73]. Corporation [Yam74]. Corresponding [Siv65]. Coup [Yaj62]. Crane [Nar76]. created [Ein82]. crossroads [Kan74]. Cultural [SH78, Wat70b]. culture [Mat78, Aok71]. Cumulative [Ano71a]. Current [Mur76]. Cyclic [Nak64b]. Cyclical [Siv65]. Cycloid [Har71]. Cycloïde [Har71].

D’Alembert [Yam71]. Dales [Tak76]. dans [Har69, Koy78]. Darwin [Sib62]. Data [Ito66]. Day [SS66, Sib62]. Days [II62]. Debate [Tom65]. Debation [Tom64]. Decade [Shi65]. Decagon [YL71]. Definition [Tom65]. delaying [Tan63]. Deo [Aok71]. Descartes [Nak74, Yam68]. descriptions [Ots75]. dessen [Tan72]. Determinations [Nak63]. Dettonville [Har69]. d’Euclide [Mur75b]. deux [Yaj63a]. Developed [Wat62]. Development [AN78, Got77, Has64, Hir62a, Ish62, Nak62b, Nak66, Nis67, Tak79, Tan68, Tan69, Tan77b, Ter76, Yag67, Yag72, Sai62, Mur71]. Diana [Nar76]. Did [Nak62b, Nak74]. Diffusion [Kob64, Nar76]. Dijksterhuis [Taj74, Wat66b]. Discourses [Yam77a]. Discovery [Nak68, Wat78]. Disk [Mur68]. Dispersion [Yag67]. Dissemination [Kob64]. Distributions [TO78]. Documents [Pre74]. d’oeil [Yaj62]. Dr. [Ish70]. d’un [Yaj71]. during [HN65, yH77, Nak65e, Shi65, Tan65, Tan77a, Wat64, vdP75]. Dutch [II62]. Dynamic [Wat62]. Dynamics [Wat66b]. Dynamism [Aok71]. Dynasties [Yab63].

E. [Wat66b]. Early [Bla73, Nak65e, Nis69, War69, Dôk73, SW67, Wat76]. Ecological [Sib62]. ed [Eza77, Mur78, TH70, Wat74, Yaj68]. Edited [Ish73, Ish74, Ish75, Yua74]. Education [FY68]. Educational [Nak66]. Edward [Wat74]. Effect [ES67]. Einführung [Tan67]. Einige [Ots75, Tan67, Tan72]. Einstein [Eza77, Eza77, Eza79, Nis79, Oga79].

Electric [Ima78]. électricité [Yam76, Yam77b]. electricity [Yam76, Yam77b]. Electrodynamics [Hir66]. Electromagnetic [Hir62a]. Electron [ES67]. Electrons [Hir62a]. electrostatic [Yam77b]. électrostatique [Yam77b]. Éléments [Mur75b]. Éloge [Kan74]. Emphasis [SW67]. emptiness [Koy78]. Encyclopedical [Mur78]. Encyclopedists [Kaw76a]. Endocrinology [NGD66]. Engi [TO78]. Engi-Shiki [TO78]. Engine [Tak79]. engineer [Yaj71]. Engineers [Ima78]. Enrichment [Mac75]. Entretien [Yaj63b]. Entwicklung [Sai62]. Environmental [SH78]. Equations [Lev70]. Era [Mur78]. ère [Yaj68]. Errata [Yag69]. ersten [Tan76]. Erythraean [SS66]. Especially [Taj75, Tom62]. Essays [Aok71]. Essential [Tan64]. Established [Ish74]. Establishment [Dok69]. état [Mur76]. Ether [Hir68]. Ethnographical [Mac75]. Euclid [Ito66]. Euclide [Mur75b]. eugenics [Suz75]. Europe [Mac78]. European [Kob64]. Européennes [Kob64]. Evaluation [Sib62]. Evidence [Oga79]. Evolution [Sin69, Tan69, WO68]. Evolutionist [Wat66a]. ex [Aok71]. expected [Shi72]. experience [Koy78, Yam76, Yam76, Koy78]. Experiment [Oga79, Yok72b]. Experiments [Iwa77]. Explanations [Yua74]. explanatory [Ish75]. Explosives [Nam70]. expressed [Siv65]. Externalist [Nak72].

F [Aok73]. Factors [Ish62]. Facts [Mur75a]. Fall [HN68]. Fang [YL72]. Fellmann [Aok73]. Ferdinand [Aok73]. Feudal [Tan77a]. Feudalistic [Dôk73]. Few [HN65, Mur67]. Field [Hir62a]. Fields [HN68]. Figures [Har69]. final [Yam76]. finale [Yam76]. Fire [Iwa77]. Fire-making [Iwa77]. First [Ach66, Ita63, Tan76]. Fishery [Suz74]. Force [Taj75, Taj76].

Forerunner [Kob76]. Forgotten [YL72]. Formalization [iT75]. Formation [HN64, Mic79, Nak62c, Nis73]. Forschung [Tan62]. Forschungen [Tan65]. found [BM72, Tom62]. Founder [Kaw76b]. francais [Yaj71]. Franklin [Yam76, Yam77b]. French [Har69, Har71, Kaw76a, Kob64, Koy78, Mur76, Yaj62, Yaj63a, Yaj63b, Yaj71, Yam68, Yam71, Yam76, Yam77b, Yos77]. Frühgeschichte [Hir70b]. Fundamental [Ish62, Sai62]. future [Shi72].

Gōryū [Nak68]. Gakkai [Shi69].

Gakujutsu [Ish73]. Gakushiin [Ish73]. Galileo [Nak62c, Yam77a]. gegen [Tan66]. Gegenwart [Kan74]. gegenwärtigen [Tan72]. General [Tak67, WO68]. Genesis [HN70]. Geneticists [Suz75]. Genetics [Sin69, Yas64]. Genron [Mur75b]. Genryu [Miu79]. Geographical [TO78]. Geology [Pre74]. Geophysical [Kim72]. German [Kan74, Ots75, Sai62, Tan62, Tan63, Tan64, Tan65, Tan66, Tan67, Tan72, Tan76]. Geschichte [Aok73, Tan62, Tan63, Tan66, Tan72, Tan76]. Giichi [Ano77a]. Gijutsu [TH70, TH72, TH73]. Gijutsushi [Yam73]. Gillispie [Shi66]. glance [Yaj62]. Glass [Mur68]. God [Nak64a]. goiter [Ots75]. Grant [Wat74]. Gravenhage [Mac75]. Group [Hir70a, Nak70, Yag71]. Growth [Yua70]. Grundbegriffe [Tan76]. Grunde [Tan63]. Grundproblem [Sai62]. Guenpaku [Yaj63b]. Gulick [Wat66a]. Gyōseki [Ish74].

H [Nak67a]. H. [ES67]. Hakubutsugaku [Kim74]. Hakubutsugaku-shi [Kim74]. Hales [Yag66]. Half [Nak65e]. Hamai [Kob76]. Hanaoka [Tak67]. Hantaro [Yag71]. Hara [Tak77b]. Harvard [Wat74, Yab69]. Harvey [Nak65c]. Hattenshi [Got77]. Hawkins [Mur71].

Heading [TH73]. Health [Has64]. Heart [Nak77]. Heat [Wat62]. Heibon [Kim74]. Heibon-sha [Kim74]. Heinrich [Ish70]. Henry [Mat75]. her [Tan77a]. Hermann [Hir70b]. Herschel [Wat70a, Wat78]. heterogeneous [Yaj63a]. hétérogènes [Yaj63a]. Hideo [Yua74]. Hidetaka [Mur75b]. Hirayama [Yua74]. Hirose [Yua74]. Hirosige [Eza77, Kan74, Nak76, Ano74a, Kan74]. Hiroto [Mur78]. Histoire [Kob64, Yaj68, Mur76, Yaj62, Yam71]. Historians [HN65, Nak72, Shi65, Suz72]. Historical [Bla74, Got77, Ish62, Kon78, SH78, Kan74, Wat70a]. historische [Kan74]. History [Ano69c, Ano78e, Ano79c, Ano79d, AN78, Has64, Hir62b, HN65, Hir70a, Hsi68, Ish73, Ish74, Ish78, Kam63, Kim74, Kob64, Miu62, Mur66, Mur76, Nak62a, Nak65a, Nak65b, Nak76, Nak62d, Nak65e, Nak69, Nak70, Nem74, ÔD73, iO62, Pre74, SW67, Sat79b, Shi69, Suz72, Tak77b, iT69, Tan69, Ter76, TH70, TH75, TY65, TY68, Wat71, WP85, Yab69, Yag75, Yam62, YT62, Yas64, Yos62b, Yos62a, Yua71, Ots77, Tan62, Tan63, Tan66, Tan72, Tan76, Yaj62, Yam71, TH70, Yam74]. Honda [Kaw76b]. Hora [Ish73]. Hoyanagi [Ish74]. humaine [Yaj63a]. human [Yaj63a]. Humanism [Shi66]. Hundert [Tan64, Tan65]. Hundred [Tan64, Tan65]. Hundredth [Yag75]. Hutton [Wat78]. Hypothese [Tan63]. Hypotheses [Ait75]. Hypothesis [Hir71, Tan66, Tan63].

iaponicus [Yaj73]. Ibid [Wat70a]. ibn [Lev71]. Ichi [Yam73, Yam73, Yam74]. Idea [Tan77b]. Igakushiso [Kaw79]. II [Tan63, Tan65, Bla74, FY68, Ohm64, Tan63, Tan65, Yag72]. Iida [Yam73, Yam74]. III [Har71, Tan66, Tan66]. Ikeda [Mur75b]. Image [iT75]. Impact [Yok72b, Yab69]. Importation [Mac78, Tom64]. Inō [Ish74, Ish74]. including [Ish75]. incorporation [Tan67]. Independent [Nak68, Tan65]. Indeterminate [Lev70]. Index [Ano71a, TH72]. Indian [Oha67]. induction [Yam77b]. Industries [Tom62]. inertia [Yaj63a]. inertie [Yaj63a]. Influence [Kaw76a, Yag66]. Information [Ano70a, Ano71b, Ano72a, Ano73a, Ano74b, Ano76a, Ano77c, Ano78b, Ano79a]. Infrared [Wat78]. ingénieur [Yaj71]. Innovation [Tom65]. Insita [Taj75]. Institutions [Nak66]. Instruments [Mac74]. Integration [Mur71]. Interest [Nak62a]. Interpretation [Kon78, Taj74]. Introduction [Mac74, Mac77, Nak64a, Tan67]. Invented [Nam70]. Investigations [Wat62]. Invisible [Nar76]. Iran [Mur70]. Iron [Yam73]. irrégularité [Har69]. irregularity [Har69]. Isotope [Tan77b]. Istanbul [Lev70]. Itô [Mur75b]. Ito [Miu79, Tak77a, Tak77b]. IV [Tan69]. IX [Ano71a]. Izumi [Mur68].

J [Wat66b]. Jahre [Tan64]. Jahren [Tan65]. Jahrhunderts [Tan66]. Jai [Bla74]. James [Wat78]. Jan [Mac77]. Japan [Ano79d, Ish73, Ish74, Ish75, Kim74, Mat78, Nak76, Shi69, Tan64, Tan65, Tan67, TH70, Yam73, Ach66, Ano78e, AN78, Bar76, Dok69, Dôk73, Fuj75, Has64, Hig75, Hir62b, Hir63, Hir70a, Ima78, II62, Kam63, Kaw76b, Kob76, Kob64, Mac74, Mac77, Mic79, Miu62, Miy77, Mur76, Nag70, Nak62a, Nak65a, Nak65b, Nak64a, Nak64b, Nem74, Nis79, ÔD73, Ots77, Pre74, Sai62, Sat79b, Sib62, SS66, Suz68, Suz75, iT69, Tan64, Tan65, Tan67, Tan68, Tan76, Tan77a, Tan77b, Ter76, Tom64, Tom65, TH75, TY65, TY68, Wat64, Wat66a, WO68, Wat71, Yag75, Yaj62, Yaj71, Yam62, Yam74, YT62, Yua63, Yua70, Yua71]. Japanese [Mat78, Mur78, Yab69, Ano69c, HN65, yH77, Ita63, Kim77, Mac75, Mac78, Mic77, Mur68, Nak62d, Nak65e, Nak72, Oga79, iO62, Shi65, Suz72, Wat70b, Yos62b, Yos62a, vdP75, Ano79c]. Japans [Tan64, Tan76]. Japon [Yaj68, Kob64, Mur76, Yaj62, Yaj71]. japonais [Yaj68]. Jingōki [Shi77]. Jiro [Kob76]. Johannes [Ait75]. John [Wat70a, War69, Wat66a]. Joseph [Wat70a, Nak67b]. Joule [Wat62]. Journal [Ano79d, Ano78e]. Jr [Aok71]. Jutsu [Mic77].

Kāmil [YL71]. Kaempfer [Kim76]. Kagaku [Mat78, Mur78, TH70, TH72, TH73, Nak76, Tak77a]. Kagaku-Gijutsu-Shi [TH72, TH73, TH70]. Kagakushi [Yaj68]. Kagakusi [Ano78e, Ano79d]. Kagakuteki [Ish74]. Kaiso [Tan76, Tan76]. Kamil [Lev70]. Kamo [Ano77a]. Kanzō [Suz74]. Karanī [Hay77]. Karanī-Operation [Hay77]. Karel [Mac77]. Kazuo [Yua74]. Kazutaka [Ish75]. Ken [Yam73, Yam74]. Ken-Ichi [Yam73, Yam74]. Kenkyu [Ano79d, Eza77, Ano78e]. Kepler [Ait75, Nak68]. Key [TH73, TH75]. Key-Words [TH75, TH73]. Ki [Sai62]. Kikaigijutsushi [Ish73]. Kikuchi [Ish73]. Kindaikagaku [Mat78, Miu79]. klassischen [Tan72]. Knowledge [Nar76, Oga79, Tan65]. Known [Miy69]. Kochizu [Ish75]. Kodansha [Ish75]. Kokon [Ish74]. Konyō [Sip72]. Korea [SW67]. Koshiro [Mur75b]. kosmologische [Aok73]. Kotaro [Kaw76b]. Koten [Mur78]. Kropf [Ots75]. Kyôritsu [Mur75b]. Kyôritsu-Shuppan [Mur75b]. Kyo [Mur68]. Kyo-zuka [Mur68]. Kyoiku [Yua74]. Kyoshi [Mat78]. Kyoto [Mur68, Nak70]. Kyozuka [Mur68].

labor [Shi72]. Land [Ish74]. Land-Survey [Ish74]. Landmark [Ish74]. Langevin [Cuv71]. Last [HN65, yH77, Shi65]. Late [WO68, Wat70b]. Latin [Ito66]. Law [Nak68, Taj74]. Laws [Ita63]. Learning [II62]. Lebesgue [Mur71]. Leiden [Yam76, vdP75]. Length [Nak63]. letters [Har69, Yaj71]. Lettres [Har69, Yaj71]. Levy [TO78]. Leyde [Yam76]. Librairie [Mur75b]. Licensing [Nak67a]. Liebig [Dok67]. Light [Wat78, Yag67]. Limnology [Ish78]. List [TH73]. Lorentz [ES67, Hir62a]. Lower [Nak77]. LXX [SS66]. Lynn [Aok71].

Machina [Aok71]. Machine [Tom62]. Machine-Tool [Tom62]. Magic [Wat64]. maintenance [Yaj63b]. making [Iwa77]. man [Shi72]. Mandrake [Miy66]. Manifested [Ish62]. Mansur [Lev71]. Manufacture [Tom62]. Manufacturings [Ish76]. Manuscript [Lev70]. Maps [Ish75]. Market [Suz68]. Marlik [BM72, Mur70]. Masao [Mat78, Mat78]. Mass [Aok71]. Massachusetts [Wat74]. Masuzo [Ish78, Kim74]. Materiae [WY70]. Materials [SS66]. Materienbegriffe [Tan67]. Mathematical [Mic79]. Mathematician [Ita63]. Mathematics [yH77, Mic77, Mur67, Mur76, iO62, Shi65, Shi69, Tak77b, vdP75]. mathématiques [Mur76]. Matter [Yag72, Tan67]. Meaning [Mur66]. Means [TH72]. Measurement [Sat79a]. Mechanical [Ish73]. Mechanics [Eza79]. Mediaeval [NGD66]. Medical [Kaw79, Nak67a]. Medicinals [SS66]. Medicine [Nak62a, Nak65a, Ots75]. Medieval [Wat74, Ito66, Lev71]. Medizin [Ots75]. Meiji [Ish73, Mat78, Wat76, Yaj68, Wat64, Ish73, Yaj68]. Meiji-Restoration [Ish73]. Meiji-zen [Ish73, Yaj68]. Memorial [Ano78a, Eza77]. Memoriam [Ano74a, Ano77a, Ano77b]. Mendel [Sin69]. Mental [Siv65]. Metals [Kaw76b]. Meteorology [Nem74]. Method [Siv65]. Methodological [Yas64, Tan72]. methodologischen [Tan72]. Michelson [Oga79]. Mid [Ish75]. Mid-19th [Ish75]. Middle [Tak76, Miy69]. Milk [Suz65, Suz68]. Milne [War69]. Minoru [Ano78a]. Minting [Lev71]. Mirai [Shi72].

Mirror [Wat64]. Missed [Kaw76a]. Missionary [Wat66a]. Mitteilung [Tan64, Tan65]. Mittelalters [Aok73]. Miyamoto [Kaw79]. Miye [Mur75b]. Model [HN70, Nis67, Yag64, Yag67, Yag69, Yag72]. Models [Tan69]. Modern [Hir69, Mat78, Mic79, Miu79, Nak76, Nak63, Tak67, Tan64, Yam62, Yos62b, Tan62, Tan67]. Modernen [Tan62, Tan64, Tan67]. Modernization [Mur78]. Monotani [Mur68]. Mori [Kaw79, Kaw79]. Morley [Oga79]. Mosbach [Hir70b]. Mound [Mur68]. Movement [Nak77, Suz75]. Movements [Tom62]. Münster [Aok73]. Murata [Tak77b]. Muroga [Ish75]. Museums [AN78]. Mutsumi [Ish74].

N [Aok73]. Nagaoka [Yag69, Yag71, Kim72, Yag64, Yag67, Yag72, YK73]. Nagatsune [Ish76]. Nakamura [Mur75b]. Nakaoka [Shi72]. Nakayama [Yab69]. Natural [Kim74, Mur75a, Sai62]. Nature [Yam77a]. Naturphilosophie [Sai62]. Naturwissenschaften [Tan64]. Needham [Nak67b]. Netherlands [Mac78]. neunzehnten [Tan66]. News [Ano69a, Ano70b, Ano71c, Ano72b, Ano73b, Ano74c, Ano76b, Ano77d, Ano78c, Ano79b]. Newton [Nak62c, Taj74, Taj75, Taj76, Wat66b, WY70]. Newtonian [Yag66]. ni [Tak77a]. Night [Yaj63b]. Nihon [Shi69, Kim74, TH70, TH73]. Nihonjin [Mat78]. Nikon [TH72]. Nineteenth [WO68, Wat70b, Tan66, Yua63]. Ningen [Shi72]. Nippon [Ish73, Yaj68, Yam74, Mur78, Yam73, Ish75]. Nisio [Eza77]. No [Ano69c, Ano74d, Bar76, Yag69, Got77, Ish74, Kaw79, Miu79, Nak76, Shi72]. Nobuo [Ish75]. nocturne [Yaj63b]. Nollet [Yam76]. Nomenclature [Kim77]. Northern [Pre74]. Nos [Ano79c]. Note [iO62, Tan68]. Notes [Bla73]. Notion [Tan69]. Nuclear [Hir63].

Nucleus [Nis65]. numbering [Har69].

Numérotation [Har69].

Obscure [Wat78]. Oda [Ish75]. Ogai [Kaw79]. Oil [Ish76]. okeru [Tak77a]. Okura [Ish76]. Old [Mic79, Mur68, Nis69, Sai62]. ombre [Yaj63b]. Once [Miy66]. One [Tan64, Tan65]. only [Miu62, Nak62b, Tom62, Yos62b]. Operation [Hay77]. Opinions [Taj74, Tan66, Tom62]. Organic [Nak67b]. Organization [Kam63]. Origin [BM72, Ita63, Tom64, Yok72b]. Origins [Hir65, Mur71, Tan66]. Osaka [Yua74, Mur68]. Outline [TH73]. Oyatoi [Mat78].

p [Ish75]. Pacific [Mat78, Wat70b].

Pancreas [Miy69]. Papers

[Eza77, Mic79, Sin69]. Parameters [Nak64b]. Part [Tan64, Tan64, Tan65]. Pascal [Har71, Koy78]. Pasteur [Dok67]. Patent [Tom64]. Patents [Tom64]. Paul [Cuv71]. Pendulums [Yok72b]. Pennsylvania [Wat70a]. pensée [Yaj63a]. Pentagon [YL71]. period [Tan65, Wat64, Wat76, vdP75]. Periode [Tan65]. Periodicals [WO68]. Periplus [SS66]. Perplexities [Tom62]. personality [Kan74]. Persönlichkeit [Kan74]. phase [Yam76]. Philosopher [Nak67b]. Philosophie [Aok73]. Philosophy [Nag70, Wat70a, Sai62]. Physical [Kim72, Tan69, Wat70a, Yag75]. Physics [Hir62b, Hir63, HN65, HN68, Hir69, Yam68, WP85, Yam76, Hir70a]. Physik [Hir70b]. Physiology [Nak65c, Nak77]. Physique [Yam68, Yam76]. Pi [Mur68]. Pioneer [Dôk73, Tak67]. Planck [HN70]. Plant [Kim77]. Plants [Mac78, TO78]. Policy [Sat79b]. Political [Tom62]. pp [Ish74, Mur75b, Wat70a, Yua74, Aok71, Hir70b, Shi72]. pp.

[Kim74, Mur71, Wat74, Yam73]. Pre [Mur78, Nak63, Hir63]. Pre-Modern [Nak63]. Pre-Modernization [Mur78]. Pre-War [Hir63]. Precision [Kob76]. Predating [Wat78]. Prehistory [Tan67]. Present [Sib62, SS66, Kan74, Tan72]. Present-day [Sib62]. Presentations [Mic79]. Preservation [Tan64]. Press [Aok71, Mur71, Wat70a, Wat74, Yab69]. Principle [Nis66]. Probabilistic [Kon78]. probabilités [Yam71]. probabilities [Yam71]. Problem [Nak62c, Sai62]. Probleme [Tan67, Tan72]. Problems [SH78, Shi77, Tan67, Yas64, Mat78, Tan72]. Process [Tan64]. Processes [Mic79]. Production [Tom62]. Products [Suz65]. Program [Bla74]. Progress [Tan64, Shi72]. Promotion [Ish73]. Proposition [Sat79a]. Proto [NGD66]. Proto-Endocrinology [NGD66]. Prozess [Tan64]. Public [Has64]. Publication [Hig75, Yag71]. Published [Ish73, Ish74, Ish75, Nak65b, Pre74].

Quantentheorie [Hir70b]. Quantitas [WY70]. Quantitative [TH75]. Quantum [Eza79, Got77, Nis69, Nis73]. Quasi [TH75]. Quasi-Quantitative [TH75]. Quelques [Yam71]. qu’on [Yaj63a].

R [Nak67a]. Radiation [HN70, Wat78]. Radioactivity [Tan77b]. Raja [Bla74]. Rangaku [Kat65, Sip72]. Rarities [Mac75]. Rays [Nis65, Nis69]. reasons [Tan63]. Recently [Nak65b]. Reception [Tan77a, Tan77b, Tan76]. recherches [Mur76]. recognition [Tan63]. Reconstruction [Sat79a, Siv67]. Recreative [Shi77]. Reexamined [Wat66b]. Reference [Nak77, iT75]. Reform [Aok73]. Rekishô [Ohm63, Ohm64]. Relation [SS66, Tan69]. Relativity [Cuv71, Hir65, Hir66, Hir68, Ein82]. Remarks [Mur67]. rencontre [Yaj63a]. Report [Iwa77]. Research [Kam63, Mur76, YK73, Tan62, Tan65, Yag71]. Researches [Hir63, Kim72]. Resemblance [Mic79]. Responses [Mat78]. Restoration [Ish73]. Review [Aok71, Aok73, AN78, Eza77, Got77, Hir70b, Ish73, Ish74, Ish75, Ish76, Ish78, Kaw79, Kim74, Mat78, Miu79, Mur78, Mur71, Mur75b, Nak67a, Nak76, Nar76, Sat79b, Shi72, Tak76, Tak77a, Tak77b, Tom65, TH70, Wat70a, Wat74, Yab69, Yaj68, Yam73, Yam74, Yua74]. Revival [Nak64b]. Revolution [Bar76, Yua63, Mat75]. Rezeption [Tan76]. Richard [Tak76]. Rikusuigakushi [Ish78]. Rise [HN68]. Rodo [Shi72]. Roku [Ish76]. Role [Kim72, Mur75a, Nis67, Tan62]. Rolle [Tan62]. Ron [Yam73]. Roots [Kon78, Miu79]. Royal [Mac75]. Ryokichi [Tak77b]. Ryoshiron [Got77].

’s [Mac75]. ’s-Gravenhage [Mac75]. Saigusa [Mur78]. Sakoku [vdP75]. Samples [Mur70]. San [Yam73]. San-Ichi [Yam73]. Saturnian [Yag69, Yag64, Yag67, Yag72]. Sawai [Bla74]. Scheidewege [Kan74]. Scholars [Kat65, Yos62a]. Scholastik [Aok73]. Schools [Kat65]. Science [Ano69c, Ano78e, Ano79d, AN78, FY68, Hsi68, Ish73, Kaw76b, Kob64, Mat78, Miu79, Mur78, Nag70, Nak76, Nak65d, Nak69, Nak71, Nak72, SW67, Sat79b, Shi66, Tak77a, Tan64, TH70, TH75, Wat70b, Wat71, Wat76, Yok72a, Yos62a, Yua71, Yaj62, Nak70, Wat74, Ano79c]. Sciences [Kob64, Mur75a, Wat70a, Yaj62, Yaj68]. Scientific [Bar76, Bla73, Ish74, Mac74, Nak66, Nar76, Tak76, War69, Wat70b, Yua63, Yua70]. Scientist [Dôk73, Mat75, Suz74]. Scientists [SW67]. Sea [SS66]. Second [Taj74]. see [Yaj63a]. Seen [Lev70]. Seimi [Tan76, Tan76]. seinem [Tan76]. Seishu [Tak67]. Seiyu [Ish76]. Seizo [Ano77b]. Sekaizu [Ish75]. Sekaizu-hen [Ish75]. Seki [Yua74]. Selbstädigung [Tan65]. Selbständigung [Tan64]. Self [Tan64]. Self-Preservation [Tan64]. Septuagint [SS66]. Sequence [Ish62]. Series [Mur78]. Set [Mur66]. Shôjô [SS66]. Shôjô-In [SS66]. sha [Kim74]. shadow [Yaj63b]. Shakaishi [Nak76]. Shi [TH72, TH73, TH70, Kim74]. Shift [Yua62]. Shigeru [Yab69]. Shiki [TO78]. Shimodaira [Yua74]. Shinkokai [Ish73]. Shinobu [Kaw79]. Shinsho [Ohm63, Ohm64]. Shobo [Yam73]. SHoin [Ish74]. Short [Miu62, iO62, SS66, Ots77]. Shryock [Nak67a]. Shuntaro [Mur75b, Miu79, Tak77a, Tak77b]. Shuppan [Mur75b]. Siebold [Kim76, Mac78]. Sigeko [Eza77]. Significance [Mac77]. Silicosis [Miu62]. Simple [Siv65]. Singh [Bla74]. Sir [Wat70a]. Skeptical [Tan66]. skeptischer [Tan66]. Social [Has64, Hir63, Nak76]. Society [Bla73, Hig75, Ish73, Shi69, TH70, Yag75]. Soddy [Tan77b]. Some [Mic77, Mur70, Nem74, Taj74, Tom62, TH75, Yam71, Ots75, Tan67, Tan72]. Sommerfeld [Nis73]. Source [Hig75, Hir69, Mur78, Wat74]. Special [Nak77]. Species [Kim77]. Specimens [Mac75]. Spectroscopy [YK73]. SS [Aok73]. Stage [Nis69]. Stages [Hsi68, Ish62]. Standpoint [Sin69]. State [Mur76]. States [Pre74]. Statistical [Eza79, WO68, YK73]. Steam [Tak79]. Steel [Yam73, Yam74]. Stephan [Yag66]. Stoney [iT75]. Structure [Nis69, Yag72, Tan72]. Struktur [Tan72]. Students [vdP75]. Studied [Wat64]. Studien [Tan64]. Studies [Ano69c, Ano79c, Hir62b, HN65, II62, Kim72, Mur68, Nak65b, Nak62d, ÔD73, Suz72, Tan64, Tan77b, TY65, Wat70a, Yos62b, Yos62a]. Study [Bla73, Mur70, Ohm63, Ohm64, Ter76, YT62]. Substance [Tan69]. Success [Wat70a]. Sugakushi [Shi69, Tak77b]. Suggestions [Ano69b]. Suguita [Yaj63b]. Sujet [Har71].

Support [Nak74]. Surgery [Tak67].

Survey

[yH77, Ish74, Nak62a, Nak65a, Nak71]. Sutra [Mur68]. Synopsis [Nak69]. System [Tom64, Yok72a].

T. [Nak77]. Table [Ano69c, Ano70c,

Ano72c, Ano73c, Ano74d, Ano76c, Ano77e, Ano78d, Ano78e, Ano79c, Ano79d]. Tables [Yab63]. Tablets [Mic79]. Tadataka [Ish74]. Taikei [TH70, TH72, TH73]. Taisei [Ish75]. Takabayashi [Got77]. Takakazu [Yua74]. Takehiko [Got77]. Takeo [Ish75]. Tamotsu [Tak77b]. Tanaka [Ano78a]. Taste [TY68]. Teachers [Mat78, Wat76]. Teaching [iT69, Wat71]. Technics [Ish62, Tom65]. Techniques [Iwa77]. technological [Shi72]. Technology [Hsi68, Ish73, Kob76, Mac77, Mur78, TH70, Yam62, Yam73, Yos62b, Yos62a, Yua71]. Teil [Tan64]. Tekkô [Yam73]. Tepe [BM72]. Terasaka [Mur75b]. term [Tan72]. Terms [Tan76]. Tetsuro [Shi72]. Tetu [Eza77, Kan74, Ano74a, Kan74, Nak76]. Their [Kim72, SH78]. Them [Mic77]. Theologie [Aok73]. théorie [Yam76, Yos77]. Theorien [Sai62]. Theories [Nak65c, Sib62, Sai62]. Theory [Cuv71, ES67, Got77, Hir62a, HN64, Hir65, Hir66, Hir68, HN70, Mur66, Mur71, Nak74, Nak77, Nis69, Nis73, Tak79, Tan77a, iT73, iT75, Wat62, Wat66b, WO68, Yag72, Ein82, Yam76, Yos77]. There [Bar76]. Thesaurus [TH72, TH73]. Thinking [Miy77, Mur67]. Third [Nak68]. Thomas [Wat66a, Mur71]. those [Mic79]. Thought [Kaw79, Nak66, Yaj63a]. Thumberg [Kim76]. Tokugawa [Bar76]. Tokyo [Ish73, Ish74, Ish75, Ish78, Kim74, Mur75b, Yaj68]. Tomb [Mur68]. Tomio [Ish73]. Tool [Tom62]. Toshiyoshi [Ish73]. Tosho [Yua74]. Total [TH72]. Traditional [Mic77, Yos62b]. traduits [Mur75b]. Trajectory [II62]. transition [Tan72].

Translation [Ito66]. Transmission [Lev70, Nis79]. Transplanting [Tan64]. Traveled [Miy66]. Treatise [Lev71]. Trend [WO68]. Trepidation [Nak64b]. Trial [TH75]. Tropical [FY68, Nak63]. Ts’ui [YL72]. Turn [HN68]. Two [Sin69, Yaj63a].

Übergang [Tan72]. Uchimura [Suz74]. Udagawa [Tan76, Dôk73, Tan76]. Ueno [Ish78, Kim74]. Understanding [SW67]. United [Pre74]. Universities [FY68, Ter76]. University [Mur71, Wat70a, Wat74, Yab69, vdP75]. Un’no [Ish75]. until [Ish75]. Ursprünge [Tan66]. Used [TO78, Wat62].

V [Tan72, Tan72]. Variation [Nak64b]. Various [HN68]. Verlag [Hir70b]. Verpflanzung [Tan64]. Verspätung [Tan63]. via [Mac78]. vide [Koy78]. Viewed [Sin69]. Viewpoint [SH78]. vii [Ish78]. Virtualité [Mur66]. Vis [Taj75]. void [Koy78]. voit [Yaj63a]. Vol [Wat70a]. Vols [TH70]. Volume [Eza77, Ish75]. Volumes [Ano71a, Yaj68]. vom [Sai62]. Vorgeschichte [Tan67].

während [Tan65]. Wallis [Har71]. War [FY68, Hir63]. Was [Bar76]. Wasan [Kaw76a]. Watanabe [Mat78]. Way [Mur67]. Weapon [Tom62]. Weise [Tan65]. Weite [Aok71]. Werdegang [Tan62]. Werdegangs [Tan64]. Werk [Tan76]. wesentlicher [Tan64]. West [Nak66]. Western [Aok71, Mat78, Yab69, Mac77, Siv65]. Westernization [Tom62]. Who [Nam70, Ish74]. William [Nak65c]. Willis [Nak77]. Wisconsin [Mur71]. without [Ait75]. Words [TH75, TH73]. Work [Bla73, Nis79, War69, Yag66, Tan76]. Works [Mic77, Yua74]. World [Ish75, Nak62c, FY68, Suz65]. World-Formation [Nak62c]. Write [Yag75]. Written [Ish76]. Wurtz [Yos77]. Wutai [Yab63].

X [Nis69]. X-rays [Nis69]. xii [Ish74]. xviii [Wat74].

Yôan [Dôk73]. Year [Nak63, Siv65]. Years [HN65, Tan64, Tan65]. Yoan [Tan76, Tan76].

Zeeman [ES67]. zen [Ish73, Yaj68]. Zensho [Mur78]. zuka [Mur68]. zum [Tan72]. zur [Aok73, Kan74, Tan62, Tan63, Tan66, Tan72, Tan76].

References

Achiwa:1966:FAC

[Ach66] Goro Achiwa. On the first anatomical chart in Japan. Japanese Studies in the History of Science, 5 (??):193–206, ???? 1966. CODEN JSHIAE. ISSN 0090-0176.

Aiton:1975:JKA

[Ait75] Eric Aiton. Johannes Kepler and the astronomy without hypotheses. Japanese Studies in the History of Science, 14(??):49–71, ???? 1975. CODEN JSHIAE. ISSN 0090-0176.

Aoki:1978:RHD

[AN78] Kunio Aoki and Tohru Nakagawa. A review of the history of the development of science museums in Japan. Japanese Studies in the History of Science, 17(??):1–12, ???? 1978. CODEN JSHIAE. ISSN 0090-0176.

Anonymous:1969:N

[Ano69a] Anonymous. News. Japanese Studies in the History of Science, 8(??):

176, ???? 1969. CODEN JSHIAE. ISSN 0090-0176.

Anonymous:1969:SC

[Ano69b] Anonymous. Suggestions for contributors. Japanese Studies in the History of Science, 8(??):185–186, ???? 1969. CODEN JSHIAE. ISSN 0090-0176.

Anonymous:1969:TCJ

[Ano69c] Anonymous. Table of contents:

Japanese Studies in the History of Science, No. 1/1962–No. 7/1968. Japanese Studies in the History of Science, 8(??):179–184, ???? 1969. CODEN JSHIAE. ISSN 0090-0176.

Anonymous:1970:IC

[Ano70a] Anonymous. Information for contributors. Japanese Studies in the History of Science, 9(??):177–178, ???? 1970. CODEN JSHIAE. ISSN 0090-0176.

Anonymous:1970:N

[Ano70b] Anonymous. News. Japanese Studies in the History of Science, 9 (??):168–169, ???? 1970. CODEN JSHIAE. ISSN 0090-0176.

Anonymous:1970:TC

[Ano70c] Anonymous. Table of contents.

Japanese Studies in the History of Science, 9(??):170–176, ???? 1970. CODEN JSHIAE. ISSN 0090-0176.

Anonymous:1971:CIV

[Ano71a] Anonymous. Cumulative index to volumes I–IX. Japanese Studies in the History of Science, 10(??):170– 186, ???? 1971. CODEN JSHIAE. ISSN 0090-0176.

Anonymous:1971:IC

[Ano71b] Anonymous. Information for contributors. Japanese Studies in the History of Science, 10(??):187–188, ???? 1971. CODEN JSHIAE. ISSN 0090-0176.

Anonymous:1971:N

[Ano71c] Anonymous. News. Japanese Studies in the History of Science, 10 (??):168–169, ???? 1971. CODEN JSHIAE. ISSN 0090-0176.

Anonymous:1972:IC

[Ano72a] Anonymous. Information for contributors. Japanese Studies in the History of Science, 11(??):171–172, ???? 1972. CODEN JSHIAE. ISSN 0090-0176.

Anonymous:1972:N

[Ano72b] Anonymous. News. Japanese Studies in the History of Science, 11 (??):166–169, ???? 1972. CODEN JSHIAE. ISSN 0090-0176.

Anonymous:1972:TC

[Ano72c] Anonymous. Table of contents. Japanese Studies in the History of Science, 11(??):170, ???? 1972. CODEN JSHIAE. ISSN 0090-0176.

Anonymous:1973:IC

[Ano73a] Anonymous. Information for contributors. Japanese Studies in the History of Science, 12(??): 155–156, ???? 1973. CO-DEN JSHIAE. ISSN 0090-0176.

                                        URL http://www.isc.meiji.ac.
                                        jp/~sano/hssj-open/Table_of_
                                        Contents-JSHS-No_12-p154.pdf.

Anonymous:1973:N

[Ano73b] Anonymous. News. Japanese Studies in the History of Science, 12 (??):152–153, ???? 1973. CO-DEN JSHIAE. ISSN 0090-0176.

URL http://www.isc.meiji.ac.
jp/~sano/hssj-open/News-JSHS-
No_12-pp152-153.pdf.

Anonymous:1973:TC

[Ano73c] Anonymous. Table of contents.

Japanese Studies in the History of Science, 12(??):154, ???? 1973. CODEN JSHIAE. ISSN 0090-0176.

URL http://www.isc.meiji.ac.
jp/~sano/hssj-open/Table_of_
Contents-JSHS-No_12-p154.pdf.

Anonymous:1974:MTH

[Ano74a] Anonymous. In memoriam: Tetu Hirosige (1928–1975). Japanese Studies in the History of Science, 13(??):148, ???? 1974. CODEN JSHIAE. ISSN 0090-0176.

Anonymous:1974:IC

[Ano74b] Anonymous. Information for contributors. Japanese Studies in the History of Science, 13(??):161–162, ???? 1974. CODEN JSHIAE. ISSN 0090-0176.

Anonymous:1974:N

[Ano74c] Anonymous. News. Japanese Studies in the History of Science, 13 (??):156–159, ???? 1974. CODEN JSHIAE. ISSN 0090-0176.

Anonymous:1974:TCN

[Ano74d] Anonymous. Table of contents no.

12 (1973). Japanese Studies in the History of Science, 13(??):160, ????

1974. CODEN JSHIAE. ISSN 0090-0176.

Anonymous:1976:IC

[Ano76a] Anonymous. Information for contributors. Japanese Studies in the History of Science, 15(??):181–182, ???? 1976. CODEN JSHIAE. ISSN 0090-0176.

Anonymous:1976:N

[Ano76b] Anonymous. News. Japanese Studies in the History of Science, 15 (??):175–178, ???? 1976. CODEN JSHIAE. ISSN 0090-0176.

Anonymous:1976:TC

[Ano76c] Anonymous. Table of contents. Japanese Studies in the History of Science, 15(??):179–180, ???? 1976. CODEN JSHIAE. ISSN 0090-0176.

Anonymous:1977:MGK

[Ano77a] Anonymous. In memoriam: Giichi Kamo 1899–1977. Japanese Studies in the History of Science, 16(??): 154, ???? 1977. CODEN JSHIAE. ISSN 0090-0176.

Anonymous:1977:MSA

[Ano77b] Anonymous. In memoriam: Seizo Aoki 1926–1977. Japanese Studies in the History of Science, 16(??): 155, ???? 1977. CODEN JSHIAE. ISSN 0090-0176.

Anonymous:1977:IC

[Ano77c] Anonymous. Information for contributors. Japanese Studies in the History of Science, 16(??):165–166, ???? 1977. CODEN JSHIAE. ISSN 0090-0176.

Anonymous:1977:N

[Ano77d] Anonymous. News. Japanese Studies in the History of Science, 16 (??):156–162, ???? 1977. CODEN JSHIAE. ISSN 0090-0176.

Anonymous:1977:TC

[Ano77e] Anonymous. Table of contents.

Japanese Studies in the History of Science, 16(??):163–164, ???? 1977. CODEN JSHIAE. ISSN 0090-0176.

Anonymous:1978:MMT

[Ano78a] Anonymous. In memorial: Minoru Tanaka. Japanese Studies in the History of Science, 17(??):158, ???? 1978. CODEN JSHIAE. ISSN 0090-0176.

Anonymous:1978:IC

[Ano78b] Anonymous. Information for contributors. Japanese Studies in the History of Science, 17(??):165–166, ???? 1978. CODEN JSHIAE. ISSN 0090-0176.

Anonymous:1978:N

[Ano78c] Anonymous. News. Japanese Studies in the History of Science, 17 (??):159–161, ???? 1978. CODEN JSHIAE. ISSN 0090-0176.

Anonymous:1978:TC

[Ano78d] Anonymous. Table of contents.

Japanese Studies in the History of Science, 17(??):162, ???? 1978. CODEN JSHIAE. ISSN 0090-0176.

Anonymous:1978:TCK

[Ano78e] Anonymous. Table of contents of Kagakusi Kenkyu (Journal of History of Science, Japan). Japanese Studies in the History of Science, 17 (??):163–164, ???? 1978. CODEN JSHIAE. ISSN 0090-0176.

Anonymous:1979:IC

[Ano79a] Anonymous. Information for contributors. Japanese Studies in the History of Science, 18(??):146–147, ???? 1979. CODEN JSHIAE. ISSN 0090-0176.

Anonymous:1979:N

[Ano79b] Anonymous. News. Japanese Studies in the History of Science, 18 (??):133–135, ???? 1979. CODEN JSHIAE. ISSN 0090-0176.

Anonymous:1979:TCJ

[Ano79c] Anonymous. Table of contents of Japanese Studies in the History of Science (nos. 10–17). Japanese Studies in the History of Science, 18 (??):136–143, ???? 1979. CODEN JSHIAE. ISSN 0090-0176.

Anonymous:1979:TCK

[Ano79d] Anonymous. Table of contents of Kagakusi Kenkyu (Journal of History of Science, Japan). Japanese Studies in the History of Science, 18 (??):144–145, ???? 1979. CODEN JSHIAE. ISSN 0090-0176.

Aoki:1971:BRL

[Aok71] Seizo Aoki. Book review: Lynn Weite, Jr.: Machina ex Deo: Essays in the Dynamism of Western Culture, Cambridge, Mass., The MIT Press, 1968, 186 pp. Japanese Studies in the History of Science, 10 (??):163–164, ???? 1971. CODEN JSHIAE. ISSN 0090-0176.

Aoki:1973:BRF

[Aok73] Seizo Aoki. Book review: Ferdinand Fellmann: Scholastik und kosmologische Reform (“Beiträge zur Geschichte der Philosophie und Theologie des Mittelalters,” N. F. Bd. 6) Münster, Aschendorff, 1971, SS. 70. Japanese Studies in the History of Science, 12 (??):145–146, ???? 1973. CO-DEN JSHIAE. ISSN 0090-0176.

URL http://www.isc.meiji.ac.
jp/~sano/hssj-open/Aoki_S-
1974-Ferdinand_Fellmann-JSHS-
No_12-pp145-146.pdf.

Bartholomew:1976:WWT

[Bar76] James R. Bartholomew. Why was there no scientific revolution in Tokugawa Japan? Japanese Studies in the History of Science, 15 (??):111–125, ???? 1976. CODEN JSHIAE. ISSN 0090-0176.

Blanpied:1973:NSE

[Bla73] William A. Blanpied. Notes for a study on the early scientific work of the Asiatic Society of Bengal. Japanese Studies in the History of Science, 12(??):121–144, ???? 1973. CODEN JSHIAE. ISSN 0090-0176.

URL http://www.isc.meiji.ac.
jp/~sano/hssj/pdf/Blanpied_W-
1974-Asiatic_Society_of_
Bengal-JSHS-No_12-pp121-144.
pdf.

Blanpied:1974:APR

[Bla74] William A. Blanpied. The astronomical program of Raja Sawai Jai Singh II and its historical context. Japanese Studies in the History of Science, 13(??):87–126, ???? 1974. CODEN JSHIAE. ISSN 0090-0176.

Beck:1972:OAF

[BM72] Cult W. Beck and Teruko Muroga. The origin of the amber found at Tepe Marlik. Japanese Studies in the History of Science, 11(??):99– 102, ???? 1972. CODEN JSHIAE. ISSN 0090-0176.

Cuvaj:1971:PLT

[Cuv71] Camillo Cuvaj. Paul Langevin and the theory of relativity. Japanese Studies in the History of Science, 10 (??):113–142, ???? 1971. CODEN JSHIAE. ISSN 0090-0176.

Doke:1967:CBL

[Dok67] Tatsumasa Doke. The controversy between Liebig and Pasteur. Japanese Studies in the History of Science, 6(??):87–95, ???? 1967. CODEN JSHIAE. ISSN 0090-0176.

Doke:1969:EBJ

[Dok69] Tatsumasa Doke. Establishment of biochemistry in Japan. Japanese Studies in the History of Science, 8 (??):145–153, ???? 1969. CODEN JSHIAE. ISSN 0090-0176.

Doke:1973:YUP

[Dôk73] Tatsumasa Dôke. Yôan Udagawa — a pioneer scientist of Early 19th Century Feudalistic Japan. Japanese Studies in the History of Science, 12(??):99–120, ???? 1973. CODEN JSHIAE. ISSN 0090-0176.

                                        URL http://www.isc.meiji.ac.
                                        jp/~sano/hssj/pdf/Doke_T-
                                        1974-Yoan   Udagawa-JSHS-No_
                                        12-pp99-120.pdf.

Einstein:1982:HCT

[Ein82] Albert Einstein. How I created the theory of relativity. Physics Today, 35(8):45–47, August 1982. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http:

//www.physicstoday.org/
resource/1/phtoad/

v35/i8/p45_s1. Introduction and translation by Yoshimasa A. Ono of a lecture given by Einstein in Kyoto, Japan on 14 December 1922. It was first published in the Japanese periodical Kaizo in 1923, and more than a half-century later, in a partial English translation in [Oga79]. Reprinted in [WP85, 243–245].

Endo:1967:ZET

[ES67] Shinji Endo and Sachie Saito. Zeeman effect and the theory of electron of H. A. Lorentz. Japanese Studies in the History of Science, 6(??):1–18, ???? 1967. CODEN JSHIAE. ISSN 0090-0176.

Ezawa:1977:BRS

[Eza77] Hiroshi Ezawa. Book review:

Sigeko Nisio (ed.): Einstein Kenkyu (Papers on Einstein: A Memorial Volume for Tetu Hirosige). Japanese Studies in the History of Science, 16(??):146–149, ???? 1977. CODEN JSHIAE. ISSN 0090-0176.

Ezawa:1979:ECSb

[Eza79] Hiroshi Ezawa. Einstein’s contributions to statistical mechanics, classical and quantum. Japanese Studies in the History of Science, 18 (??):27–72, ???? 1979. CODEN JSHIAE. ISSN 0090-0176.

Fujii:1975:AJ

[Fuj75] Kiyohisa Fujii. Atomism in Japan, 1868–1888. Japanese Studies in the History of Science, 14(??):141–156, ???? 1975. CODEN JSHIAE. ISSN 0090-0176.

Fujita:1968:SET

[FY68] Chie Fujita and Eri Yagi. Science education in tropical African universities after World War II. Japanese Studies in the History of Science, 7(??):143–158, ???? 1968. CODEN JSHIAE. ISSN 0090-0176.

Goto:1977:BRT

[Got77] Kunio Gotō. Book review: Takehiko Takabayashi: Ryoshiron no Hattenshi (The Historical Development of Quantum Theory). Japanese Studies in the History of Science, 16(??):150–153, ???? 1977. CODEN JSHIAE. ISSN 0090-0176.

Hara:1969:INF

[Har69] Kokiti Hara. Sur l’irrégularité de numérotation des figures dans les lettres de Dettonville. (French) [On the irregularity of numbering figures in the letters of Dettonville]. Japanese Studies in the History of Science, 8(??):33–54, ???? 1969. CODEN JSHIAE. ISSN 0090-0176.

Hara:1971:PWA

[Har71] Kokiti Hara. Pascal et Wallis au sujet de la cycloïde (iii). (French) [Pascal and Wallis on the cycloid (iii)]. Japanese Studies in the History of Science, 10(??):95–112, ???? 1971. CODEN JSHIAE. ISSN 0090-0176.

Hashimoto:1964:DSC

[Has64] Michio Hashimoto. Development of social consciousness in the history of public health in Japan. Japanese Studies in the History of Science, 3 (??):127–??, ???? 1964. CODEN JSHIAE. ISSN 0090-0176.

Hayashi:1977:KKO

[Hay77] Takao Hayashi. Karan.ī in the Karan.ī-operation. Japanese Studies in the History of Science, 16 (??):51–59, ???? 1977. CODEN JSHIAE. ISSN 0090-0176.

Higasi:1975:RAC

[Hig75] Keniti Higasi. Recent activities of the Chemical Society of Japan and the publication of a source book in chemistry. Japanese Studies in the History of Science, 14(??):1–10, ???? 1975. CODEN JSHIAE. ISSN 0090-0176.

Hirosige:1962:LTE

[Hir62a] Tetu Hirosige. Lorentz’s theory of electrons and the development of the concept of electromagnetic field. Japanese Studies in the History of Science, 1(??):101–110, ???? 1962. CODEN JSHIAE. ISSN 0090-0176.

Hirosige:1962:SHP

[Hir62b] Tetu Hirosige. Studies of history of physics in Japan. Japanese Studies in the History of Science, 1(??):26– 34, ???? 1962. CODEN JSHIAE. ISSN 0090-0176.

Hirosige:1963:SCR

[Hir63] Tetu Hirosige. Social conditions for the researches of nuclear physics in Pre-War Japan. Japanese Studies in the History of Science, 2(??):80– 93, ???? 1963. CODEN JSHIAE. ISSN 0090-0176.

Hirosige:1965:CCO

[Hir65] Tetu Hirosige. A consideration concerning the origins of the theory of relativity. Japanese Studies in the History of Science, 4(??):117–123, ???? 1965. CODEN JSHIAE. ISSN 0090-0176.

Hirosige:1966:EBT

[Hir66] Tetu Hirosige. Electrodynamics before the theory of relativity, 1890– 1905. Japanese Studies in the History of Science, 5(??):1–49, ???? 1966. CODEN JSHIAE. ISSN 0090-0176.

Hirosige:1968:TRE

[Hir68] Tetu Hirosige. Theory of relativity and the ether. Japanese Studies in the History of Science, 7(??):37–54, ???? 1968. CODEN JSHIAE. ISSN 0090-0176.

Hirosige:1969:SBM

[Hir69] Tetu Hirosige. Source books in the modern physics. Japanese Studies in the History of Science, 8(??):17– 20, ???? 1969. CODEN JSHIAE. ISSN 0090-0176.

Hirosige:1970:AJG

[Hir70a] Tetu Hirosige. Activities of Japan’s Group for History of Physics. Japanese Studies in the History of Science, 9(??):5–12, ???? 1970. CODEN JSHIAE. ISSN 0090-0176.

Hirosige:1970:BRA

[Hir70b] Tetu Hirosige. Book review: Armin Hermann: Frühgeschichte der Quantentheorie (1899–1913), Physik Verlag, Mosbach in Baden, 1969, 181 pp. Japanese Studies in the History of Science, 9(??):161– 163, ???? 1970. CODEN JSHIAE. ISSN 0090-0176.

Hirosige:1971:VBH

[Hir71] Tetu Hirosige. The van den Broek hypothesis. Japanese Studies in the History of Science, 10(??):143–162, ???? 1971. CODEN JSHIAE. ISSN 0090-0176.

Hirosige:1964:FBT

[HN64] Tetu Hirosige and Sigeko Nisio. Formation of Bohr’s theory of atomic constitution. Japanese Studies in the History of Science, 3(??):6–28, ???? 1964. CODEN JSHIAE. ISSN 0090-0176.

Hirosige:1965:SHP

[HN65] Tetu Hirosige and Sigeko Nisio.

Studies in the history of physics by Japanese historians during the last few years. Japanese Studies in the History of Science, 4(??):28–40, ???? 1965. CODEN JSHIAE. ISSN 0090-0176.

Hirosige:1968:RFV

[HN68] Tetu Hirosige and Sigeko Nisio.

Rise and fall of various fields of physics at the turn of the century. Japanese Studies in the History of Science, 7(??):93–114, ???? 1968. CODEN JSHIAE. ISSN 0090-0176.

Hirosige:1970:GBA

[HN70] Tetu Hirosige and Sigeko Nisio. The genesis of the Bohr atom model and Planck’s theory of radiation. Japanese Studies in the History of Science, 9(??):35–47, ???? 1970. CODEN JSHIAE. ISSN 0090-0176.

Hsieh:1968:SHC

[Hsi68] Shin-Hui Hsieh. Stages in the history of contemporary science and technology. Japanese Studies in the History of Science, 7(??):115–128, ???? 1968. CODEN JSHIAE. ISSN 0090-0176.

Itakura:1962:STJ

[II62]  Kiyonobu  Itakura

and Reiko Itakura. Studies of trajectory in Japan before the days of Dutch learning. Japanese Studies in the History of Science, 1(??):83–93, ???? 1962. CODEN JSHIAE. ISSN 0090-0176.

Imazu:1978:BEE

[Ima78] Kenji Imazu. The beginning of electric engineers in Japan. Japanese Studies in the History of Science, 17(??):13–26, ???? 1978. CODEN JSHIAE. ISSN 0090-0176.

Ova:1962:SNH

[iO62] Shin ichi Ova. A short note on the history of Japanese mathematics. Japanese Studies in the History of Science, 1(??):23–25, ???? 1962. CODEN JSHIAE. ISSN 0090-0176.

Ishigai:1962:FFD

[Ish62] Seikan Ishigai. Fundamental factors in the development of technics as manifested in the sequence of its historical stages. Japanese Studies in the History of Science, 1 (??):125–134, ???? 1962. CODEN JSHIAE. ISSN 0090-0176.

Ishiyama:1970:BCD

[Ish70] Yoshikazu Ishiyama. A brief chronology of Dr. Heinrich Burger. Japanese Studies in the History of Science, 9(??):107–113, ???? 1970. CODEN JSHIAE. ISSN 0090-0176.

Ishiyama:1973:BRM

[Ish73] Hiroshi Ishiyama. Book review: Meiji-zen Nippon Kikaigijutsushi (History of Mechanical Technology in Japan before Meiji-Restoration) By Tomio Hora and Toshiyoshi Kikuchi: Edited by Nippon Gakushiin (Japan Academy). Published by Nippon Gakujutsu Shinkokai (Japan Society for Promotion of Science), Tokyo, 1973. 16 + 351 + 42pp. Japanese Studies in the History of Science, 12 (??):149–151, ???? 1973. CO-DEN JSHIAE. ISSN 0090-0176.

URL http://www.isc.meiji.ac.
jp/~sano/hssj-open/Ishiyama_
H-1974-Meiji-zen_Nippon_
Kikaigijutsushi-JSHS-No_12-
pp149-151.pdf.

Ishiyama:1974:BRI

[Ish74] Hiroshi Ishiyama. Book review: Inō Tadataka no Kagakuteki Gyōseki (A New Appreciation of the Scientific Achievement of Inō Tadataka, who Established a Landmark in the History of Land-Survey and Cartography of Japan). Edited by Mutsumi Hoyanagi. Published by Kokon SHoin, Tokyo, 1974. xii + 510 pp, Yen 8,800. Japanese Studies in the History of Science, 13 (??):143–145, ???? 1974. CODEN JSHIAE. ISSN 0090-0176.

Ishiyama:1975:BRP

[Ish75] Hiroshi Ishiyama. Book review:

Nippon Kochizu Taisei: Sekaizuhen (World in Japan Maps until the Mid-19th Century) Edited by Takeo Oda, Nobuo Muroga, and Kazutaka Un’no. Published by Kodansha, Tokyo, 1975. 284 p. (including 138 maps) with explanatory volume 96 p. 43.5 × 31 cm. Yen 62,000. Japanese Studies in the History of Science, 14(??):168–170, ???? 1975. CODEN JSHIAE. ISSN 0090-0176.

Ishiyama:1976:BRS

[Ish76] Hiroshi Ishiyama. Book review: Seiyu Roku: On Oil Manufacturings. Written by Okura Nagatsune. Japanese Studies in the History of Science, 15(??):173–174, ???? 1976. CODEN JSHIAE. ISSN 0090-0176.

Ishiyama:1978:BRM

[Ish78] Hiroshi Ishiyama. Book review: Masuzo Ueno: Rikusuigakushi (The History of Limnology), Baifûkan, Tokyo, 1977. vii + 367. Japanese Studies in the History of Science, 17 (??):150–153, ???? 1978. CODEN JSHIAE. ISSN 0090-0176.

Tamamushi:1969:THC

[iT69] Bun ichi Tamamushi. Teaching the history of chemistry in Japan. Japanese Studies in the History of Science, 8(??):9–16, ???? 1969. CODEN JSHIAE. ISSN 0090-0176.

Tsuneishi:1973:AT

[iT73] Kei ichi Tsuneishi. On the Abbe theory (1873). Japanese Studies in the History of Science, 12 (??):79–91, ???? 1973. CO-DEN JSHIAE. ISSN 0090-0176.

                                        URL http://www.isc.meiji.ac.
                                        jp/~sano/hssj/pdf/Tsuneishi_
                                        K-1974-Abbe_Theory-JSHS-No_
                                        12-pp79-91.pdf.

Tsuneishi:1975:SCI

[iT75] Kei ichi Tsuneishi. On Stoney’s concept of image, with reference to the formalization of the Abbe theory. Japanese Studies in the History of Science, 14(??):95–102, ???? 1975. CODEN JSHIAE. ISSN 0090-0176.

Itakura:1963:FBL

[Ita63] Kiyonobu Itakura. The first ballistical laws by a Japanese mathematician and its origin. Japanese Studies in the History of Science, 2 (??):136–145, ???? 1963. CODEN JSHIAE. ISSN 0090-0176.

Ito:1966:MLT

[Ito66] Shuntaro Ito. On the Medieval Latin translation of the data of Euclid. Japanese Studies in the History of Science, 5(??):100–118, ???? 1966. CODEN JSHIAE. ISSN 0090-0176.

Iwaki:1977:REA

[Iwa77] Masao Iwaki. A report of experiments on ancient fire-making techniques. Japanese Studies in the History of Science, 16(??):91–93, ???? 1977. CODEN JSHIAE. ISSN 0090-0176.

Kamatani:1963:HRO

[Kam63] Chikayoshi Kamatani. The history of research organization in Japan. Japanese Studies in the History of Science, 2(??):1–79, ???? 1963. CODEN JSHIAE. ISSN 0090-0176.

Kangro:1974:ETH

[Kan74] Hans Kangro. Éloge: Tetu Hirosige:

historische Persönlichkeit am Scheidewege zur Gegenwart. (German)

[Éloge: Tetu Hirosige: historical personality at the crossroads of the present]. Japanese Studies in the History of Science, 13(??):149–155, ???? 1974. CODEN JSHIAE. ISSN 0090-0176.

Katagiri:1965:SRS

[Kat65] Kazuo Katagiri. The schools of Rangaku scholars. Japanese Studies in the History of Science, 4 (??):173–??, ???? 1965. CODEN JSHIAE. ISSN 0090-0176.

Kawajiri:1976:MIF

[Kaw76a] Nobuo Kawajiri. The missed influence of the French Encyclopedists on Wasan. Japanese Studies in the History of Science, 15(??):79– 95, ???? 1976. CODEN JSHIAE. ISSN 0090-0176.

Kawamiya:1976:KHF

[Kaw76b] Nobuo Kawamiya. Kotaro Honda: Founder of the science of metals in Japan. Japanese Studies in the History of Science, 15(??):145–158, ???? 1976. CODEN JSHIAE. ISSN 0090-0176.

Kawamoto:1979:BRS

[Kaw79] Hideo Kawamoto. Book review: Shinobu Miyamoto: Mori Ogai no Igakushiso (The Medical Thought of Ogai Mori). Japanese Studies in the History of Science, 18(??):127– 129, ???? 1979. CODEN JSHIAE. ISSN 0090-0176.

Kimura:1972:NGS

[Kim72] Tosaku Kimura. Nagaoka’s geophysical studies and their role in his physical researches. Japanese Studies in the History of Science, 11(??):91–98, ???? 1972. CODEN JSHIAE. ISSN 0090-0176.

Kimura:1974:BRM

[Kim74] Yojiro Kimura. Book review: Masuzo Ueno: Nihon Hakubutsugakushi (History of Natural History in Japan), Heibon-sha, Tokyo, 1973, 680 + 72 pp. Japanese Studies in the History of Science, 13(??):141– 142, ???? 1974. CODEN JSHIAE. ISSN 0090-0176.

Kimura:1976:KTS

[Kim76] Yojiro Kimura. Kaempfer, Thumberg and Siebold. Japanese Studies in the History of Science, 15(??):1– 13, ???? 1976. CODEN JSHIAE. ISSN 0090-0176.

Kimura:1977:BNJ

[Kim77] Yojiro Kimura. Binominal nomenclature of the Japanese plant species. Japanese Studies in the History of Science, 16(??):105–108, ???? 1977. CODEN JSHIAE. ISSN 0090-0176.

Kobori:1964:AHD

[Kob64] Akira Kobori. Un aspect de l’histoire de la diffusion des sciences européennes au Japon. (French) [An aspect of the history of the dissemination of European science in Japan]. Japanese Studies in the History of Science, 3(??):1–5, ???? 1964. CODEN JSHIAE. ISSN 0090-0176.

Kobayashi:1976:BBJ

[Kob76] Tatsuya Kobayashi. Brief biography of Jiro Hamai — a forerunner of precision technology in Japan. Japanese Studies in the History of Science, 15(??):159–162, ???? 1976. CODEN JSHIAE. ISSN 0090-0176.

Konno:1978:HRB

[Kon78] Hiroyuki Konno. The historical roots of Born’s probabilistic interpretation. Japanese Studies in the History of Science, 17(??):129–145, ???? 1978. CODEN JSHIAE. ISSN 0090-0176.

Koyanagi:1978:PED

[Koy78] Kimiyo Koyanagi. Pascal et l’Éxperience du vide dans le vide. (French) [Pascal and the experience of emptiness in the void]. Japanese Studies in the History of Science, 17 (??):105–127, ???? 1978. CODEN JSHIAE. ISSN 0090-0176.

Levey:1970:TIE

[Lev70] Martin Levey. Transmission of indeterminate equations as seen in an Istanbul manuscript of Abu Kamil. Japanese Studies in the History of Science, 9(??):17–25, ???? 1970. CODEN JSHIAE. ISSN 0090-0176.

Levey:1971:CAM

[Lev71] Martin Levey. Chemical aspects of Medieval Arabic minting in a treatise by Mansur ibn Bacra. Japanese Studies in the History of Science, 10(S1):1–136, ???? 1971. CODEN JSHIAE. ISSN 0090-

                                        0176. URL http://hssj.info/ja/
                                        archives/641.

MacLean:1974:IBS

[Mac74] J. MacLean. The introduction of books and scientific instruments into Japan. Japanese Studies in the History of Science, 13(??):9–68, ???? 1974. CODEN JSHIAE. ISSN 0090-0176.

MacLean:1975:ERC

[Mac75] J. MacLean. The enrichment of the Royal Cabinet of Rarities at ’s-Gravenhage with Japanese ethnographical specimens from 1815 to 1848. Japanese Studies in the History of Science, 14(??):117–139, ???? 1975. CODEN JSHIAE. ISSN 0090-0176.

MacLean:1977:SJK

[Mac77] J. MacLean. The significance of Jan Karel van den Broek (1814–1865) for the introduction of Western technology into Japan. Japanese Studies in the History of Science, 16(??):69–90, ???? 1977. CODEN JSHIAE. ISSN 0090-0176.

MacLean:1978:SIJ

[Mac78] J. MacLean. Von Siebold and the importation of Japanese plants into Europe via The Netherlands. Japanese Studies in the History of Science, 17(??):43–79, ???? 1978. CODEN JSHIAE. ISSN 0090-0176.

Matsuo:1975:HCS

[Mat75] Yukitoshi Matsuo. Henry Cavendish: A scientist in the age of the Révolution Chimique. Japanese Studies in the History of Science, 14(??):83–94, ???? 1975. CODEN JSHIAE. ISSN 0090-0176.

Matsuo:1978:BRM

[Mat78] Yukitoshi Matsuo. Book review:

Masao Watanabe: Oyatoi Beikokujin Kagaku Kyoshi (Science Across the Pacific — American Science Teachers in Meiji Japan). Masao Watanabe: Nihonjin to Kindaikagaku (Modern Science and the Japanese — Responses to the Western culture and the contemporary problems). Japanese Studies in the History of Science, 17(??):154–157, ???? 1978. CODEN JSHIAE. ISSN 0090-0176.

Michiwaki:1977:SJT

[Mic77] Yoshimasa Michiwaki. On some jutsu in traditional Japanese mathematics and some recent works about them. Japanese Studies in the History of Science, 16(??):61– 68, ???? 1977. CODEN JSHIAE. ISSN 0090-0176.

Michiwaki:1979:RBP

[Mic79] Yoshimasa Michiwaki. On the resemblance between the presentations and the formation processes of modern mathematical papers and those of old mathematical tablets in Japan. Japanese Studies in the History of Science, 18(??):117–125, ???? 1979. CODEN JSHIAE. ISSN 0090-0176.

Miura:1962:SHS

[Miu62] Toyohiko Miura. A short history of silicosis in Japan [abstract only]. Japanese Studies in the History of Science, 1(??):137, ???? 1962. CO-DEN JSHIAE. ISSN 0090-0176.

Miura:1979:BRS

[Miu79] Nobuo Miura. Book review: Shuntaro Ito: Kindaikagaku no Genryu (Roots of Modern Science). Japanese Studies in the History of Science, 18(??):130–132, ???? 1979. CODEN JSHIAE. ISSN 0090-0176.

Miyashita:1966:MOT

[Miy66] Saburō Miyashita. Mandrake once traveled in China as an anaesthetic. Japanese Studies in the History of Science, 5(??):189–192, ???? 1966. CODEN JSHIAE. ISSN 0090-0176.

Miyasita:1969:PKC

[Miy69] Saburō Miyasita. Pancreas known by the Chinese in the Middle Ages. Japanese Studies in the History of Science, 8(??):167–171, ???? 1969. CODEN JSHIAE. ISSN 0090-0176.

Miyake:1977:ATC

[Miy77] Kouji Miyake. Analytical thinking of Confucianism in Japan and China. Japanese Studies in the History of Science, 16(??):109–111, ???? 1977. CODEN JSHIAE. ISSN 0090-0176.

Murata:1966:MVH

[Mur66] Tamotsu Murata. On the meaning of ‘virtualité’ in the history of the set theory. Japanese Studies in the History of Science, 5(??):119–139, ???? 1966. CODEN JSHIAE. ISSN 0090-0176.

Murata:1967:FRA

[Mur67] Tamotsu Murata. A few remarks on the atomistic way of thinking in mathematics. Japanese Studies in the History of Science, 6(??):47–59, ???? 1967. CODEN JSHIAE. ISSN 0090-0176.

Muroga:1968:CSA

[Mur68] Teruko Muroga. Chemical studies on ancient Japanese glass beads from the Izumi Kyo-zuka (Sutra Mound), Osaka; the Kyozuka Old Tomb, Osaka; and the Monotani Old Tomb, Kyoto, and ancient Chinese glass disk (pi), and glass bead. Japanese Studies in the History of Science, 7(??):83–92, ???? 1968. CODEN JSHIAE. ISSN 0090-0176.

Muroga:1970:CSS

[Mur70] Teruko Muroga. A chemical study on some archaeological samples from Marlik in Iran. Japanese Studies in the History of Science, 9 (??):99–105, ???? 1970. CODEN JSHIAE. ISSN 0090-0176.

Murata:1971:BRT

[Mur71] Tamotsu Murata. Book review: Thomas Hawkins: Lebesgue’s Theory of Integration, Its Origins and Development, University of Wisconsin Press, 1970, 227 pp. Japanese Studies in the History of Science, 10 (??):165–167, ???? 1971. CODEN JSHIAE. ISSN 0090-0176.

Murakami:1975:RFN

[Mur75a] Yoichiro P. Murakami. Role of facts in natural sciences. Japanese Studies in the History of Science, 14 (??):39–47, ???? 1975. CODEN JSHIAE. ISSN 0090-0176.

Murata:1975:BRE

[Mur75b] Tamotsu Murata. Book review: Euclide Genron (Éléments d’Euclide), traduits par Koshiro Nakamura, Hidetaka Terasaka, Shuntaro Itô et Miye Ikeda. Librairie Kyôritsu-Shuppan, Tokyo, 1971, pp. 560. Japanese Studies in the History of Science, 14(??):165–167, ???? 1975. CODEN JSHIAE. ISSN 0090-0176.

Murata:1976:LAR

[Mur76] Tamotsu Murata. L’état actuel des recherches en histoire des mathématiques, au Japon. (French) [The current state of research in the history of mathematics in Japan]. Japanese Studies in the History of Science, 15(??):21–36, ???? 1976. CODEN JSHIAE. ISSN 0090-0176.

Murakami:1978:BRH

[Mur78] Yoichiro P. Murakami. Book review: Hiroto Saigusa (ed.): Nippon Kagaku Koten Zensho (Encyclopedical Source Book Series of Japanese Science and Technology in Pre-Modernization Era). Japanese Studies in the History of Science, 17 (??):147–149, ???? 1978. CODEN JSHIAE. ISSN 0090-0176.

Nagai:1970:PSJ

[Nag70] Hiroshi Nagai. Philosophy of science in Japan: 1966–1970. Japanese Studies in the History of Science, 9(??):13–16, ???? 1970. CODEN JSHIAE. ISSN 0090-0176.

Nakagawa:1962:SIH

[Nak62a] Yonezo Nakagawa. A survey of the interest for the history of medicine in Japan. Japanese Studies in the History of Science, 1(??):38–45, ???? 1962. CODEN JSHIAE. ISSN 0090-0176.

Nakayama:1962:DAC

[Nak62b] Shigeru Nakayama. Did astrology contribute to the development of astronomy? [abstract only]. Japanese Studies in the History of Science, 1(??):137, ???? 1962. CODEN JSHIAE. ISSN 0090-0176.

Nakayama:1962:GNP

[Nak62c] Shigeru Nakayama. Galileo and Newton’s problem of world-formation. Japanese Studies in the History of Science, 1(??):76– 82, ???? 1962. CODEN JSHIAE. ISSN 0090-0176.

Nakayama:1962:JSH

[Nak62d] Shigeru Nakayama. Japanese studies in the history of astronomy. Japanese Studies in the History of Science, 1(??):14–22, ???? 1962. CODEN JSHIAE. ISSN 0090-0176.

Nakayama:1963:APM

[Nak63] Shigeru Nakayama. Accuracy of pre-modern determinations of tropical year length. Japanese Studies in the History of Science, 2 (??):101–118, ???? 1963. CODEN JSHIAE. ISSN 0090-0176.

Nakayama:1964:AGI

[Nak64a] Shigeru Nakayama. Abhorrence of “God” in the introduction of Copernicanism into Japan. Japanese Studies in the History of Science, 3(??):60–67, ???? 1964. CODEN JSHIAE. ISSN 0090-0176.

Nakayama:1964:CVA

[Nak64b] Shigeru Nakayama. Cyclic variation of astronomical parameters and the revival of trepidation in Japan. Japanese Studies in the History of Science, 3(??):68–80, ???? 1964. CODEN JSHIAE. ISSN 0090-0176.

Nakagawa:1965:SHM

[Nak65a] Yonezo Nakagawa. A survey of the history of medicine in Japan, 1961– 1965. Japanese Studies in the History of Science, 4(??):76–80, ???? 1965. CODEN JSHIAE. ISSN 0090-0176.

Nakamura:1965:SHB

[Nak65b] Teiri Nakamura. Studies of the history of biology recently published in Japan. Japanese Studies in the History of Science, 4(??):60–75, ???? 1965. CODEN JSHIAE. ISSN 0090-0176.

Nakamura:1965:WHH

[Nak65c] Teiri Nakamura. William Harvey and his theories of physiology. Japanese Studies in the History of Science, 4(??):143–161, ???? 1965. CODEN JSHIAE. ISSN 0090-0176.

Nakayama:1965:CCC

[Nak65d] Shigeru Nakayama. Characteristics of Chinese calendrical science. Japanese Studies in the History of Science, 4(??):124–131, ???? 1965. CODEN JSHIAE. ISSN 0090-0176.

Nakayama:1965:JAH

[Nak65e] Shigeru Nakayama. Japanese activities in the history of astronomy during the early half of 1960’s. Japanese Studies in the History of Science, 4(??):1–19, ???? 1965. CODEN JSHIAE. ISSN 0090-0176.

Nakayama:1966:EID

[Nak66] Shigeru Nakayama. Educational institutions and the development of scientific thought in China and the West. Japanese Studies in the History of Science, 5(??):172–179, ???? 1966. CODEN JSHIAE. ISSN 0090-

Nakagawa:1967:BRM

[Nak67a] Yonezo Nakagawa. Book review: Medical Licensing in America, 1650–1965 by R. H. Shryock. Japanese Studies in the History of Science, 6(??):155–??, ???? 1967. CODEN JSHIAE. ISSN 0090-0176.

Nakayama:1967:JNO

[Nak67b] Shigeru Nakayama. Joseph Needham, organic philosopher. Japanese Studies in the History of Science, 6 (??):138–154, ???? 1967. CODEN JSHIAE. ISSN 0090-0176.

Nakayama:1968:AID

[Nak68] Shigeru Nakayama. On the alleged independent discovery of Kepler’s Third Law by Asada Gōryū. Japanese Studies in the History of Science, 7(??):55–60, ???? 1968. CODEN JSHIAE. ISSN 0090-0176.

Nakayama:1969:SHC

[Nak69] Shigeru Nakayama. Synopsis [of] the history of Chinese science. Japanese Studies in the History of Science, 8(??):21–31, ???? 1969. CODEN JSHIAE. ISSN 0090-0176.

Nakayama:1970:KGH

[Nak70] Shigeru Nakayama. Kyoto Group of the History of Chinese Science. Japanese Studies in the History of Science, 9(??):1–4, ???? 1970. CO-DEN JSHIAE. ISSN 0090-0176.

Nakayama:1971:SSC

[Nak71] Shigeru Nakayama. Survey of “science in colonialism”. Japanese Studies in the History of Science, 10(??):17–21, ???? 1971. CODEN JSHIAE. ISSN 0090-0176.

Nakayama:1972:EAJ

[Nak72] Shigeru Nakayama. Externalist approach of Japanese historians of science. Japanese Studies in the History of Science, 11(??):1–10, ???? 1972. CODEN JSHIAE. ISSN 0090-0176.

Nakamura:1974:HWD

[Nak74] Teire Nakamura. How and why did Descartes support the theory of blood circulation? Japanese Studies in the History of Science, 13 (??):75–80, ???? 1974. CODEN JSHIAE. ISSN 0090-0176.

Nakaoka:1976:BRT

[Nak76] Tetsuro Nakaoka. Book review:

Tetu Hirosige: Kagaku no Shakaishi (A Social History of Science in Modern Japan). Japanese Studies in the History of Science, 15 (??):163–168, ???? 1976. CODEN JSHIAE. ISSN 0090-0176.

Nakamura:1977:WLP

[Nak77] Teiri Nakamura. T. Willis’ and Lower’s physiology with special reference to the theory of heart movement. Japanese Studies in the History of Science, 16(??):23–41, ???? 1977. CODEN JSHIAE. ISSN 0090-0176.

Named:1970:WIE

[Nam70] Heizo Named. Who invented the explosives? Japanese Studies in the History of Science, 9(??):49–98, ???? 1970. CODEN JSHIAE. ISSN 0090-0176.

Narisada:1976:BRD

[Nar76] Kaoru Narisada. Book review: Diana Crane: Invisible Colleges; Diffusion of Knowledge in Scientific Community. Japanese Studies in the History of Science, 15(??):169– 170, ???? 1976. CODEN JSHIAE. ISSN 0090-0176.

Nemoto:1974:SCH

[Nem74] Junkichi Nemoto. Some characteristics of the history of meteorology in Japan. Japanese Studies in the History of Science, 13(??):1–8, ???? 1974. CODEN JSHIAE. ISSN 0090-0176.

Needham:1966:PEM

[NGD66] Joseph Needham and Lu Gwei-Djen. Proto-endocrinology in Mediaeval China. Japanese Studies in the History of Science, 5(??):150– 171, ???? 1966. CODEN JSHIAE. ISSN 0090-0176.

Nisio:1965:RAN

[Nis65] Sigeko Nisio. α-rays and the atomic nucleus. Japanese Studies in the History of Science, 4(??):91–116, ???? 1965. CODEN JSHIAE. ISSN 0090-0176.

Nisio:1966:BCP

[Nis66] Sigeko Nisio. From Balmer to the combination principle. Japanese Studies in the History of Science, 5(??):50–74, ???? 1966. CODEN JSHIAE. ISSN 0090-0176.

Nishio:1967:RCC

[Nis67] Sigeko Nishio. The role of the chemical considerations in the development of Bohr atom model. Japanese Studies in the History of Science, 6(??):26–40, ???? 1967. CODEN JSHIAE. ISSN 0090-0176.

Nisio:1969:XRA

[Nis69] Sigeko Nisio. X-rays and atomic structure at the early stage of the old quantum theory. Japanese Studies in the History of Science, 8 (??):55–75, ???? 1969. CODEN JSHIAE. ISSN 0090-0176.

Nisio:1973:FSQ

[Nis73] Sigeko Nisio. The formation of the Sommerfeld quantum theory of 1916. Japanese Studies in the History of Science, 12(??):39–78, ???? 1973. CODEN JSHIAE. ISSN

0090-0176. URL http://www.isc.
meiji.ac.jp/~sano/hssj/pdf/
Nisio_S-1974-Sommerfeld-JSHS-
No_12-pp39-78.pdf.

Nisio:1979:TEW

[Nis79] Sigeko Nisio. The transmission of Einstein’s work to Japan. Japanese Studies in the History of Science, 18(??):1–8, ???? 1979. CODEN JSHIAE. ISSN 0090-0176.

Onuma:1973:RSJ

[ÔD73] Masanori Ônuma and Tatsumasa Dôke. Recent studies in Japan on the history of chemistry. Japanese Studies in the History of Science, 12(??):5–14, ???? 1973. CO-DEN JSHIAE. ISSN 0090-0176.

URL http://www.isc.meiji.ac.
jp/~sano/hssj/pdf/Onuma_M-
1974-History_of_Chemistry-
JSHS-No_12-pp5-14.pdf.

Ogawa:1979:JEE

[Oga79] Tsuyoshi Ogawa. Japanese evidence for Einstein’s knowledge of the Michelson–Morley experiment. Japanese Studies in the History of Science, 18(??):73–81, ???? 1979. CODEN JSHIAE. ISSN 0090-0176. This article contains a partial translation of Einstein’s lecture How I created the theory of relativity at Kyoto University on 14 December 1922. A full translation is available in [Ein82].

Ohami:1967:IA

[Oha67] Isao Ohami. On Indian atomism (I). Japanese Studies in the History of Science, 6(??):41–46, ???? 1967. CODEN JSHIAE. ISSN 0090-0176.

Ohmori:1963:SRS

[Ohm63] Minoru Ohmori. A study on the Rekishô Shinsho. Japanese Studies in the History of Science, 2 (??):146–153, ???? 1963. CODEN JSHIAE. ISSN 0090-0176.

Ohmori:1964:SRS

[Ohm64] Minoru Ohmori. A study on the Rekishô Shinsho (II). Japanese Studies in the History of Science, 3(??):81–88, ???? 1964. CODEN JSHIAE. ISSN 0090-0176.

Otsuka:1975:EBK

[Ots75] Yasuo Otsuka. Einige Beschreibungen über Kropf in der antiken Medizin. (German) [Some descriptions about goiter in ancient medicine]. Japanese Studies in the History of Science, 14(??):103–108, ???? 1975. CODEN JSHIAE. ISSN 0090-0176.

Otsuka:1977:SHB

[Ots77] Yasuo Otsuka. A short history of bioethics in Japan. Japanese Studies in the History of Science, 16 (??):17–22, ???? 1977. CODEN JSHIAE. ISSN 0090-0176.

Prestana:1974:HGJ

[Pre74] Harold R. Prestana. History of geology in Japan and Northern Asia: Documents published by the United States Congress. Japanese Studies in the History of Science, 13 (??):69–73, ???? 1974. CODEN JSHIAE. ISSN 0090-0176.

Saigusa:1962:ETK

[Sai62] Hiroto Saigusa. Die Entwicklung der Theorien vom “Ki” (Chhi)als Grundproblem der Naturphilosophie in Alten Japan. (German) [The development of theories of “ki” (chhi) as a fundamental problem of natural philosophy in Old Japan]. Japanese Studies in the History of Science, 1(??):51–56, ???? 1962. CODEN JSHIAE. ISSN 0090-0176.

Sato:1979:AMC

[Sat79a] Tohru Sato. Archimedes’ On the Measurement of a Circle Proposition 1: An attempt at reconstruction. Japanese Studies in the History of Science, 18(??):83–99, ???? 1979. CODEN JSHIAE. ISSN 0090-0176.

Satofuka:1979:RHS

[Sat79b] Fumihiko Satofuka. The review of the history of science policy in Japan — towards an alternative policy. Japanese Studies in the History of Science, 18(??):9–25, ???? 1979. CODEN JSHIAE. ISSN 0090-0176.

Satofuka:1978:CAE

[SH78] Fumihiko Satofuka and Yukio Himiyama. Comparative approach to the environmental problems from the viewpoint of their historical, cultural backgrounds. Japanese Studies in the History of Science, 17(??):27– 42, ???? 1978. CODEN JSHIAE. ISSN 0090-0176.

Shimodaira:1965:AJH

[Shi65] Kazuo Shimodaira. Activities of Japanese historians of mathematics during the last decade. Japanese Studies in the History of Science, 4(??):20–27, ???? 1965. CODEN JSHIAE. ISSN 0090-0176.

Shimao:1966:BSH

[Shi66] Eikoh Shimao. Between science and humanism — on C. C. Gillispie. Japanese Studies in the History of Science, 5(??):211–??, ???? 1966. CODEN JSHIAE. ISSN 0090-0176.

Shimodaira:1969:AHM

[Shi69] Kazuo Shimodaira. Activities of “The History of Mathematics Society of Japan (Nihon Sugakushi Gakkai)”. Japanese Studies in the History of Science, 8(??):1–7, ???? 1969. CODEN JSHIAE. ISSN 0090-0176.

Shizume:1972:BRT

[Shi72] Yasuo Shizume. Book review: Tetsuro Nakaoka: Ningen to Rodo no Mirai (The future of man and labor — what can be expected from technological progress?) Chuokoron, 1970, 214 pp. Japanese Studies in the History of Science, 11 (??):163–165, ???? 1972. CODEN JSHIAE. ISSN 0090-0176.

Shimodaira:1977:RPJ

[Shi77] Kazuo Shimodaira. Recreative problems on “jingōki”. Japanese Studies in the History of Science, 16(??):95–103, ???? 1977. CODEN JSHIAE. ISSN 0090-0176.

Sibuya:1962:PDE

[Sib62] Kazuo Sibuya. Present-day evaluation of the ecological aspects of Darwin’s theories in Japan. Japanese Studies in the History of Science, 1 (??):117–124, ???? 1962. CODEN JSHIAE. ISSN 0090-0176.

Sinoto:1969:MTG

[Sin69] Yosito Sinotô. Mendel’s two genetics papers viewed from the standpoint of evolution. Japanese Studies in the History of Science, 8 (??):155–166, ???? 1969. CODEN JSHIAE. ISSN 0090-0176.

Sippel:1972:AKB

[Sip72] Patricia Sippel. Aoki Konyō (1698– 1769) and the beginnings of Rangaku. Japanese Studies in the History of Science, 11(??):127–162, ???? 1972. CODEN JSHIAE. ISSN 0090-0176.

Sivin:1965:SMM

[Siv65] Nathan Sivin. A simple method for mental conversion of a year expressed in cyclical characters to the corresponding year in the Western calendar. Japanese Studies in the History of Science, 4(??):132–135, ???? 1965. CODEN JSHIAE. ISSN 0090-0176.

Sivin:1967:RCA

[Siv67] Nathan Sivin. On the reconstruction of Chinese alchemy. Japanese Studies in the History of Science, 6(??):60–86, ???? 1967. CODEN JSHIAE. ISSN 0090-0176.

Suzuki:1966:SCC

[SS66] Shigeaki Suzuki and Reiko Suzuki. A short classification of commodities of the “Periplus of the Erythraean Sea” and a relation to materials in the “Septuagint” (LXX) and the “Shôjô-In” medicinals in the present day Japan. Japanese Studies in the History of Science, 5 (??):180–189, ???? 1966. CODEN JSHIAE. ISSN 0090-0176.

Suzuki:1965:MMP

[Suz65] Shigeaki Suzuki. Milk and milk products in the Ancient World. Japanese Studies in the History of Science, 4(??), 135–142 1965. CO-DEN JSHIAE. ISSN 0090-0176.

Suzuki:1968:MMJ

[Suz68] Shigeaki Suzuki. On the market milk in Japan around 1900. Japanese Studies in the History of Science, 7(??):71–82, ???? 1968. CODEN JSHIAE. ISSN 0090-0176.

Suzuki:1972:RSH

[Suz72] Zenji Suzuki. Recent studies in the history of biology by Japanese historians. Japanese Studies in the History of Science, 11(??):11–21, ???? 1972. CODEN JSHIAE. ISSN 0090-0176.

Suzuki:1974:UKF

[Suz74] Shigeaki Suzuki. Uchimura Kanzō as a fishery scientist. Japanese Studies in the History of Science, 13 (??):127–133, ???? 1974. CODEN JSHIAE. ISSN 0090-0176.

Suzuki:1975:GEM

[Suz75] Zenji Suzuki. Geneticists and the eugenics movement in Japan.

Japanese Studies in the History of Science, 14(??):157–164, ???? 1975. CODEN JSHIAE. ISSN 0090-0176.

Sang-Woon:1967:USH

[SW67] Jeon Sang-Woon. Understanding of science in history of Korea with emphasis on scientists in Early 15th Century. Japanese Studies in the History of Science, 6(??):124–137, ???? 1967. CODEN JSHIAE. ISSN 0090-0176.

Tajima:1974:SCI

[Taj74] Nobuo Tajima. Some considerations on the interpretation of Newton’s Second Law and on Dijksterhuis’ opinions. Japanese Studies in the History of Science, 13(??):81– 86, ???? 1974. CODEN JSHIAE. ISSN 0090-0176.

Tajima:1975:NCF

[Taj75] Nobuo Tajima. On Newton’s concept of force especially Vis Insita. Japanese Studies in the History of Science, 14(??):73–81, ???? 1975. CODEN JSHIAE. ISSN 0090-0176.

Tajima:1976:NCF

[Taj76] Nobuo Tajima. On Newton’s centrifugal force. Japanese Studies in the History of Science, 15(??):65– 78, ???? 1976. CODEN JSHIAE. ISSN 0090-0176.

Takebayashi:1967:SHP

[Tak67] Hiromu Takebayashi. Seishu Hanaoka “pioneer of the general anethesia and the modern surgery”. Japanese Studies in the History of Science, 6(??):115–123, ???? 1967. CODEN JSHIAE. ISSN 0090-0176.

Takahashi:1976:BRR

[Tak76] Kenichi Takahashi. Book review: Richard C. Dales: The Scientific Achievement of the Middle Ages. Japanese Studies in the History of Science, 15(??):171–172, ???? 1976. CODEN JSHIAE. ISSN 0090-0176.

Takahashi:1977:BRS

[Tak77a] Kenichi Takahashi. Book review: Shuntaro Ito: Bunmei ni okeru Kagaku (Science in Civilizations). Japanese Studies in the History of Science, 16(??):137–140, ???? 1977. CODEN JSHIAE. ISSN 0090-0176.

Takeuchi:1977:BRS

[Tak77b] Kei Takeuchi. Book review: Shuntaro Ito, Ryokichi Hara and Tamotsu Murata: Sugakushi (History of Mathematics). Japanese Studies in the History of Science, 16 (??):141–145, ???? 1977. CODEN JSHIAE. ISSN 0090-0176.

Takayama:1979:DTS

[Tak79] Susumu Takayama. Development of “the theory of the steam engine”. Japanese Studies in the History of Science, 18(??):101–116, ???? 1979. CODEN JSHIAE. ISSN 0090-0176.

Tanaka:1962:BGA

[Tan62] Minoru Tanaka. Ein Beitrag zur Geschichte der Atomistik. Über die Rolle der Chemischen Forschung beim Werdegang der Modernen Atomistik. (German) [A contribution to the history of atomism. On the role of chemical research in the history of modern atomism]. Japanese Studies in the History of Science, 1(??):111–116, ???? 1962. CODEN JSHIAE. ISSN 0090-0176.

Tanaka:1963:BGA

[Tan63] Minoru Tanaka. Ein Beitrag zur Geschichte der Atomistik (II). Über die Grunde der Verspätung der Anerkennung der Avogadroschen Hypothese. (German) [A contribution to the history of atomism (II). On the reasons for delaying the recognition of Avogadro’s hypothesis]. Japanese Studies in the History of Science, 2(??):127–135, ???? 1963. CODEN JSHIAE. ISSN 0090-0176.

Tanaka:1964:HJC

[Tan64] Minoru Tanaka. Hundert Jahre der Chemie in Japan, Studien über den Prozess der Verpflanzung und Selbständigung der Naturwissenschaften als wesentlicher Teil des Werdegangs modernen Japans (Mitteilung I). (German) [One hundred years of chemistry in Japan, studies on the process of transplanting and self-preservation of science as an essential part of the progress of modern Japan (Part I)]. Japanese Studies in the History of Science, 3(??):89–107, ???? 1964. CODEN JSHIAE. ISSN 0090-0176.

Tanaka:1965:HJC

[Tan65] Minoru Tanaka. Hundert Jahren der Chemie in Japan (Mitteilung II). Die Art und Weise der Selbstädigung chemischer Forschungen während der Periode 1901–1930. (German) [One hundred years of chemistry in Japan (Part II). The art and knowledge of independent chemical research during the period 1901–1930]. Japanese Studies in the History of Science, 4 (??):162–172, ???? 1965. CODEN JSHIAE. ISSN 0090-0176.

Tanaka:1966:USA

[Tan66] Minoru Tanaka. Über die Ursprünge skeptischer Auffassungen gegen Atomhypothese der Chemie neunzehnten Jahrhunderts. Ein Beitrag zur Geschichte der Atomistik (III). (German) [On the origins of skeptical opinions against the atomic hypothesis of Nineteenth Century chemistry. A contribution to the history of atomism (III)]. Japanese Studies in the History of Science, 5(??):87–99, ???? 1966. CODEN JSHIAE. ISSN 0090-0176.

Tanaka:1967:EPV

[Tan67] Minoru Tanaka. Einige Probleme der Vorgeschichte der Chemie in Japan. Einführung und Aufnahme der modernen Materienbegriffe. (German) [Some problems of prehistory of chemistry in Japan. Introduction and incorporation of modern concepts of matter]. Japanese Studies in the History of Science, 6(??):96–114, ???? 1967. CODEN JSHIAE. ISSN 0090-0176.

Tanaka:1968:NDC

[Tan68] Minoru Tanaka. A note on the development of chemistry in Japan. Japanese Studies in the History of Science, 7(??):61–70, ???? 1968. CODEN JSHIAE. ISSN 0090-0176.

Tanaka:1969:CPM

[Tan69] Minoru Tanaka. Chemical and physical models for atomistic notion — its conceptual development in relation to the evolution of the concept of chemical substance. A contribution to the history of atomism (IV). Japanese Studies in the History of Science, 8(??):125–143, ???? 1969. CODEN JSHIAE. ISSN 0090-0176.

Tanaka:1972:EMP

[Tan72] Minoru Tanaka. Einige methodologischen Probleme des klassischen Begriffs der chemischen Struktur und dessen Übergang zum gegenwärtigen Begriff — Beitrag zur Geschichte der Atomistik (V). (German) [Some methodological problems of the classical concept of chemical structure and its transition to the present term — contribution to the history of atomism (V)]. Japanese Studies in the History of Science, 11(??):113–126, ???? 1972. CODEN JSHIAE. ISSN 0090-0176.

Tanaka:1976:RCG

[Tan76] Minoru Tanaka. Rezeption chemischer Grundbegriffe bei dem ersten Chemiker Japans, Udagawa Yoan (1798–1846), in seinem Werk, Seimi Kaiso. Beiträge zur Geschichte der Chemie in Japan. (German) [Reception of basic chemical terms by the first chemist of Japan, Udagawa Yoan (1798–1846), in his work, Seimi Kaiso. Contributions to the history of chemistry in Japan]. Japanese Studies in the History of Science, 15(??):97–110, ???? 1976. CODEN JSHIAE. ISSN 0090-0176.

Tanaka:1977:RAT

[Tan77a] Minoru Tanaka. The reception of atomic theory in Japan during her Feudal Age. Japanese Studies in the History of Science, 16(??):113–117, ???? 1977. CODEN JSHIAE. ISSN 0090-0176.

Tanaka:1977:RSI

[Tan77b] Minoru Tanaka. The reception of Soddy’s idea of isotope and the development of radioactivity studies in Japan. Japanese Studies in the History of Science, 16(??):119–123, ???? 1977. CODEN JSHIAE. ISSN 0090-0176.

Terasaki:1976:DSH

[Ter76] Masao Terasaki. Development of the study on the history of universities in Japan. Japanese Studies in the History of Science, 15(??):15– 19, ???? 1976. CODEN JSHIAE. ISSN 0090-0176.

Tomita:1970:BRH

[TH70] Tetsuo Tomita and Kazutoshi Hattori. Book review: History of Science Society of Japan (ed.): Nihon Kagaku-Gijutsu-shi Taikei (History of Science and Technology in Japan), 25 Vols., 1964–1970. Japanese Studies in the History of Science, 9(??):164–167, ???? 1970. CODEN JSHIAE. ISSN 0090-0176.

Tomita:1972:CTT

[TH72] Tetsuo Tomita and Kazutoshi Hattori. Compilation of a thesaurus and total index for Nikon Kagaku-Gijutsu-Shi Taikei by means of a computer. Japanese Studies in the History of Science, 11(??):41–65, ???? 1972. CODEN JSHIAE. ISSN 0090-0176.

Tomita:1973:OTN

[TH73] Tetsuo Tomita and Kazutoshi Hattori. Outline of a thesaurus of Nihon Kagaku-Gijutsu-Shi Taikei with heading list of classified key-words. Japanese Studies in the History of Science, 12(??):15–38, ???? 1973. CO-DEN JSHIAE. ISSN 0090-0176.

URL http://www.isc.meiji.ac.
jp/~sano/hssj/pdf/Tomita_T-
1974-Nihon   Kagaku-Gijutsu-
Shi_Taikei-No_12-pp15-38.pdf.

Tomita:1975:SCQ

[TH75] Tetsuo Tomita and Kazutoshi Hattori. Some considerations of quasiquantitative analysis of the history of science in Japan by keywords — trial of quantitative history. Japanese Studies in the History of Science, 14(??):13–38, ???? 1975. CODEN JSHIAE. ISSN 0090-0176.

Tomita:1978:GDP

[TO78] Tetsuo Tomita and Isao Ohami.

The geographical distributions of plants used for levy in the book “Engi-Shiki”. Japanese Studies in the History of Science, 17(??):81– 96, ???? 1978. CODEN JSHIAE. ISSN 0090-0176.

Tomita:1962:SPC

[Tom62] Tetsuo Tomita. Some perplexities of the concepts of industries, found in the political opinions on the weapon manufacture, in the age of the Westernization movements in China especially on the machine-tool production [abstract only]. Japanese Studies in the History of Science, 1(??):138–??, ???? 1962. CODEN JSHIAE. ISSN 0090-

Tomita:1964:OPS

[Tom64] Tetsuo Tomita. The origin of [the] patent system in Japan — around the debation of importation patents. Japanese Studies in the History of Science, 3(??):114–126, ???? 1964. CODEN JSHIAE. ISSN 0090-0176.

Tomita:1965:IDT

[Tom65] Tetsuo Tomita. Innovation and definition of technics — a review of a recent debate in Japan. Japanese Studies in the History of Science, 4(??):81–90, ???? 1965. CODEN JSHIAE. ISSN 0090-0176.

Tsuzuki:1965:SHC

[TY65] Yojiro Tsuzuki and Aiko Yamashita. On the studies of history of chemistry in Japan. Japanese Studies in the History of Science, 4(??):41–59, ???? 1965. CODEN JSHIAE. ISSN 0090-0176.

Tsuzuki:1968:HCT

[TY68] Yojiro Tsuzuki and Aiko Yamashita. History of the chemistry of taste in Japan. Japanese Studies in the History of Science, 7(??):1– 26, ???? 1968. CODEN JSHIAE. ISSN 0090-0176.

vanderPas:1975:JSM

[vdP75] Peter W. van der Pas. Japanese students of mathematics at the University of Leiden during the Sakoku Period. Japanese Studies in the History of Science, 14(??):109–116, ???? 1975. CODEN JSHIAE. ISSN 0090-0176.

Wartnaby:1969:ESW

[War69] John Wartnaby. The early scientific work of John Milne. Japanese Studies in the History of Science, 8 (??):77–124, ???? 1969. CODEN JSHIAE. ISSN 0090-0176.

Watanabe:1962:DTH

[Wat62] Masao Watanabe. The dynamic theory of heat as used and developed by Joule in his investigations. Japanese Studies in the History of Science, 1(??):94–100, ???? 1962. CODEN JSHIAE. ISSN 0090-0176.

Watanabe:1964:MMS

[Wat64] Masao Watanabe. The magic mirror as studied in Japan during the Meiji Period. Japanese Studies in the History of Science, 3(??):48–59, ???? 1964. CODEN JSHIAE. ISSN 0090-0176.

Watanabe:1966:JTG

[Wat66a] Masao Watanabe. John Thomas Gulick: American evolutionist and missionary in Japan. Japanese Studies in the History of Science, 5 (??):140–149, ???? 1966. CODEN JSHIAE. ISSN 0090-0176.

Watanabe:1966:NTD

[Wat66b] Masao Watanabe. Newton’s theory of dynamics reexamined by E. J. Dijksterhuis. Japanese Studies in the History of Science, 5(??):207–210, ???? 1966. CODEN JSHIAE. ISSN 0090-0176.

Watanabe:1970:BRH

[Wat70a] Masao Watanabe. Book review:

Historical Studies in the Physical Sciences, Vol. 1, University of Pennsylvania Press, 1969; Joseph Agassi, “Sir John Herschel’s Philosophy of Success,” Ibid,, pp. 1–36. Japanese Studies in the History of Science, 9 (??):159–160, ???? 1970. CODEN JSHIAE. ISSN 0090-0176.

Watanabe:1970:SAP

[Wat70b] Masao Watanabe. Science across the Pacific: American–Japanese scientific and cultural contacts in the Late Nineteenth Century. Japanese Studies in the History of Science, 9(??):115–136, ???? 1970. CODEN JSHIAE. ISSN 0090-0176.

Watanabe:1971:THS

[Wat71] Masao Watanabe. Teaching history of science in Japan. Japanese Studies in the History of Science, 10 (??):25–26, ???? 1971. CODEN JSHIAE. ISSN 0090-0176.

Watanabe:1974:BRE

[Wat74] Hiroshi Watanabe. Book review: Edward Grant (ed.): A Source Book in Medieval Science, Harvard University Press, Cambridge, Massachusetts, 1974, xviii + 864 pp. Japanese Studies in the History of Science, 13(??):146–147, ???? 1974. CODEN JSHIAE. ISSN 0090-0176.

Watanabe:1976:AST

[Wat76] Masao Watanabe. American science teachers in the Early Meiji Period. Japanese Studies in the History of Science, 15(??):127–144, ???? 1976. CODEN JSHIAE. ISSN 0090-0176.

Watanabe:1978:JHO

[Wat78] Masao Watanabe. James Hutton’s “obscure light” — a discovery of infrared radiation predating Herschel’s. Japanese Studies in the History of Science, 17(??):97–104, ???? 1978. CODEN JSHIAE. ISSN 0090-0176.

Watanabe:1968:GAT

[WO68] Masao Watanabe and Yōko Ose.

General academic trend and the evolution theory in Late Nineteenth Century Japan: A statistical analysis of the contemporary periodicals. Japanese Studies in the History of Science, 7(??):129–142, ???? 1968. CODEN JSHIAE. ISSN 0090-0176.

Weart:1985:HP

[WP85] Spencer R. Weart and Melba Phillips, editors. History of physics, volume 2 of Readings from Physics Today. American Institute of Physics, Woodbury, NY, USA, 1985. ISBN 0-88318-468-0 (paperback). 375 pp. LCCN QC7 .H694 1985.

Watanabe:1970:NQM

[WY70] Masao Watanabe and Masakazu Yoshinaka. Newton’s Quantitas Materiae. Japanese Studies in the History of Science, 9(??):27–34, ???? 1970. CODEN JSHIAE. ISSN 0090-0176.

Yabuuti:1963:ATC

[Yab63] K. Yabuuti. Astronomical tables in China — from the Wutai to the Ch’ing Dynasties. Japanese Studies in the History of Science, 2(??):94– 100, ???? 1963. CODEN JSHIAE. ISSN 0090-0176.

Yabuuti:1969:BRS

[Yab69] Kiyosi Yabuuti. Book review:

Shigeru Nakayama: A History of Japanese Astronomy, Chinese Background and Western Impact, Harvard University Press, 1969. Japanese Studies in the History of Science, 8(??):173–175, ???? 1969. CODEN JSHIAE. ISSN 0090-0176.

Yagi:1964:NSA

[Yag64] Eri Yagi. On Nagaoka’s Saturnian atomic model (1903). Japanese Studies in the History of Science, 3(??):29–47, ???? 1964. CODEN JSHIAE. ISSN 0090-0176. See errata [Yag69].

Yagi:1966:SHW

[Yag66] Eri Yagi. Stephan Hales’ work in chemistry; a Newtonian influence on 18th Century chemistry. Japanese Studies in the History of Science, 5(??):75–86, ???? 1966. CODEN JSHIAE. ISSN 0090-0176.

Yagi:1967:DNS

[Yag67] Eri Yagi. The development of Nagaoka’s Saturnian atomic model I — dispersion on light — (1905). Japanese Studies in the History of Science, 6(??):19–25, ???? 1967. CODEN JSHIAE. ISSN 0090-0176.

Yagi:1969:ENS

[Yag69] Eri Yagi. Errata: On Nagaoka’s Saturnian Atomic Model (1903), No. 3 (1964). Japanese Studies in the History of Science, 8(??):177– 178, ???? 1969. CODEN JSHIAE. ISSN 0090-0176. See [Yag64].

Yagi:1971:RGC

[Yag71] Eri Yagi. Research Group of the Committee for the Publication of Hantaro Nagaoka’s Biography.

Japanese Studies in the History of Science, 10(??):23–24, ???? 1971. CODEN JSHIAE. ISSN 0090-0176.

Yagi:1972:DNS

[Yag72] Eri Yagi. The development of Nagaoka’s Saturnian atomic model II (1904–05) — Nagaoka’s theory of the structure of matter. Japanese Studies in the History of Science, 11(??):73–89, ???? 1972. CODEN JSHIAE. ISSN 0090-0176.

Yagai:1975:PSJ

[Yag75] Eri Yagai. The Physical Society of Japan to write its history on its hundredth anniversary. Japanese Studies in the History of Science, 14(??):11–12, ???? 1975. CODEN JSHIAE. ISSN 0090-0176.

Yajima:1962:COR

[Yaj62] Suketoshi Yajima. Coup d’oeil sur l’histoire des sciences au Japon. (French) [A glance at the history of science in Japan]. Japanese Studies in the History of Science, 1(??):3–6, ???? 1962. CODEN JSHIAE. ISSN 0090-0176.

Yajima:1963:IPH

[Yaj63a] S. Yajima. De l’inertie de la pensée humaine — ce qu’on voit à rencontre des deux civilisations hétérogènes. (French) [From the inertia of human thought — what we see against the two heterogeneous civilizations]. Japanese Studies in the History of Science, 2(??):154– 163, ???? 1963. CODEN JSHIAE. ISSN 0090-0176.

Yajima:1963:LNA

[Yaj63b] S. Yajima. L’entretien nocturne avec l’ombre de Suguita Guenpaku.

(French) [Night maintenance with the shadow of Suguita Guenpaku]. Japanese Studies in the History of Science, 2(??), ???? 1963. CODEN JSHIAE. ISSN 0090-0176.

Yajima:1968:BRM

[Yaj68] Suketoshi Yajima. Book review: Meiji-zen Nippon Kagakushi (Histoire des sciences en Japon avant l’ère de Meiji, en japonais), ed. Acadèmie de Japon, Tokyo, 26 volumes, 1954–1968. Japanese Studies in the History of Science, 7 (??):159–??, ???? 1968. CODEN JSHIAE. ISSN 0090-0176.

Yajima:1971:LIF

[Yaj71] Suketoshi Yajima. Lettres d’un ingénieur français en Japon de 1877 à 1881. (French) [Letters from a French engineer in Japan from 1877 to 1881]. Japanese Studies in the History of Science, 10(??):27–57, ???? 1971. CODEN JSHIAE. ISSN 0090-0176.

Yajima:1973:CI

[Yaj73] Suketoshi Yajima. Copernicus iaponicus. Japanese Studies in the History of Science, 12(??):1–3, ???? 1973. CODEN JSHIAE. ISSN

0090-0176. URL http://www.isc.
meiji.ac.jp/~sano/hssj/pdf/
Yajima_S-1974-Copernicus-
JSHS-No_12-pp1-3.pdf.

Yamazaki:1962:JCM

[Yam62] Toshio Yamazaki. Japan’s contributions to the modern history of technology. Japanese Studies in the History of Science, 1(??):45–??, ???? 1962. CODEN JSHIAE. ISSN 0090-0176.

Yamazaki:1968:PDF

[Yam68] Eizo Yamazaki. La physique de Descartes. (French) [The physics of Descartes]. Japanese Studies in the History of Science, 7(??):27–36, ???? 1968. CODEN JSHIAE. ISSN 0090-0176.

Yamazaki:1971:DCQ

[Yam71] Eizo Yamazaki. D’Alembert et Condorcet — quelques aspects de l’histoire du calcul des probabilités. (French) [D’Alembert and Condorcet — some aspects of the history of calculating probabilities]. Japanese Studies in the History of Science, 10(??):59–93, ???? 1971. CODEN JSHIAE. ISSN 0090-0176.

Yamazaki:1973:BRK

[Yam73] Toshio Yamazaki. Book review:

Ken-Ichi Iida: Nippon Tekkô Gijutsushi Ron (Technology of Iron and Steel in Japan) San-Ichi Shobo, 1973, 466 pp. Japanese Studies in the History of Science, 12 (??):147–148, ???? 1973. CO-DEN JSHIAE. ISSN 0090-0176.

URL http://www.isc.meiji.ac.
jp/~sano/hssj-open/Yamazaki_
T-1974-Ken-Ichi   Iida-JSHS-
No_12-pp147-148.pdf.

Yamazaki:1974:BRK

[Yam74] Toshio Yamazaki. Book review:

Ken-ichi Iida, History of Steel in Japan, Nippon Steel Corporation, 1973, 40pp. Japanese Studies in the History of Science, 13(??):135–136, ???? 1974. CODEN JSHIAE. ISSN 0090-0176.

Yamazaki:1976:LNB

[Yam76] Eizo Yamazaki. L’Abbé Nollet et Benjamin Franklin — une phase finale de la physique cartésienne: la théorie de la conservation de l’électricité et de l’expérience de Leyde. (French) [Abbot Nollet and Benjamin Franklin — a final phase of Cartesian physics: the theory of the conservation of electricity and the experience in Leiden]. Japanese Studies in the History of Science, 15(??):37–64, ???? 1976. CODEN JSHIAE. ISSN 0090-0176.

Yamagata:1977:GBA

[Yam77a] Yasuji Yamagata. On Galileo’s basic attitude to nature in his discourses. Japanese Studies in the History of Science, 16(??):125–127, ???? 1977. CODEN JSHIAE. ISSN 0090-0176.

Yamazaki:1977:CEI

[Yam77b] Eizo Yamazaki. La conduction de l’électricité et l’induction électrostatique avant Franklin. (French) [Conduction of electricity and electrostatic induction before Franklin]. Japanese Studies in the History of Science, 16(??):43–49, ???? 1977. CODEN JSHIAE. ISSN 0090-0176.

Yasugi:1964:MPH

[Yas64] Ryuichi Yasugi. Methodological problems in the history of genetics. Japanese Studies in the History of Science, 3(??):108–113, ???? 1964. CODEN JSHIAE. ISSN 0090-0176.

Honda:1977:SJM

[yH77] Kin ya Honda. A survey of Japanese mathematics during the last centenary. Japanese Studies in the History of Science, 16(??):1–16, ???? 1977. CODEN JSHIAE. ISSN 0090-0176.

Yagi:1973:SAN

[YK73] Eri Yagi and Tosaku Kimura. A statistical approach to Nagaoka’s research in spectroscopy. Japanese Studies in the History of Science, 12(??):93–97, ???? 1973. CO-DEN JSHIAE. ISSN 0090-0176.

URL http://www.isc.meiji.ac.
jp/~sano/hssj/pdf/Yagi_E-
1974-Nagaoka_Spectroscopy-
JSHS-No_12-pp93-97.pdf.

Yadegari:1971:AKP

[YL71] Mohammad Yadegari and Martin Levey. Abū Kāmil’s “On the Pentagon and Decagon”. Japanese Studies in the History of Science, 10(S2):1–54, ???? 1971. CODEN JSHIAE. ISSN 0090-

0176. URL http://hssj.info/ja/
archives/657.

Yoke:1972:TFF

[YL72] Ho Peng Yoke and Beda Lim. Ts’ui Fang, a forgotten 11th-century Chinese alchemist. Japanese Studies in the History of Science, 11(??):103– 112, ???? 1972. CODEN JSHIAE. ISSN 0090-0176.

Yoke:1972:SBC

[Yok72a] Ho Peng Yoke. The system of the Book of Changes and Chinese science. Japanese Studies in the History of Science, 11(??):23–39, ???? 1972. CODEN JSHIAE. ISSN 0090-

Yokoyama:1972:OEI

[Yok72b] Masahiko Yokoyama. Origin of the experiment of impact with pendulums. Japanese Studies in the History of Science, 11(??):67–72, ???? 1972. CODEN JSHIAE. ISSN 0090-0176.

Yosida:1962:SHS

[Yos62a] Mitukuni Yosida. Studies on the history of science and technology in China by Japanese scholars. Japanese Studies in the History of Science, 1(??):7–13, ???? 1962. CODEN JSHIAE. ISSN 0090-0176.

Yosida:1962:SHJ

[Yos62b] Mitukuni Yosida. Studies on the history of the Japanese traditional arts and modern technology [abstract only]. Japanese Studies in the History of Science, 1(??):135– 136, ???? 1962. CODEN JSHIAE. ISSN 0090-0176.

Yoshida:1977:CAW

[Yos77] Akira Yoshida. Charles Adolphe Wurtz et la théorie atomique. (French) [Charles Adolphe Wurtz and atomic theory]. Japanese Studies in the History of Science, 16 (??):129–135, ???? 1977. CODEN JSHIAE. ISSN 0090-0176.

Yasugi:1962:SHB

[YT62] Ryuichi Yasugi and Hisaharu Tsukuba. On the study of the history of biology in Japan. Japanese Studies in the History of Science, 1(??):35–37, ???? 1962. CODEN JSHIAE. ISSN 0090-0176.

Yuasa:1962:CAS

[Yua62] Mitutomo Yuasa. Center of activity: its shift from the 16th to the 20th century. Japanese Studies in the History of Science, 1(??):57–75, ???? 1962. CODEN JSHIAE. ISSN 0090-0176.

Yuasa:1963:SRN

[Yua63] Mitutomo Yuasa. Scientific revolution in Nineteenth Century Japan. Japanese Studies in the History of Science, 2(??):119–126, ???? 1963. CODEN JSHIAE. ISSN 0090-0176.

Yuasa:1970:GSC

[Yua70] Mitsutomo Yuasa. The growth of scientific communities in Japan. Japanese Studies in the History of Science, 9(??):137–158, ???? 1970. CODEN JSHIAE. ISSN 0090-0176.

Yuasa:1971:HST

[Yua71] Mitsutomo Yuasa. History of science and technology in Japan. Japanese Studies in the History of Science, 10(??):1–16, ???? 1971. CODEN JSHIAE. ISSN 0090-0176.

Yuasa:1974:BRA

[Yua74] Mitsutomo Yuasa. Book review: Akira Hirayama, Kazuo Shimodaira, and Hideo Hirose, Takakazu Seki’s Collected Works, edited with Explanations, Osaka Kyoiku Tosho, 1974, 872 pp. (18 × 26 cm). Japanese Studies in the History of Science, 13(??):137–140, ???? 1974. CODEN JSHIAE. ISSN 0090-0176.